### v1

# Advancing Strategic Planning and Dynamic Control of Complex Projects

Abstract
Strategic project planning and dynamic control is essential for ensuring complex projects to be executed best-fit for common purpose. When best-for-project strategies are no longer conceivable for humans, there is a need for effective and efficient computer-aided decision support. To this end, standard simulation-driven evaluation and a-posteriori decision-making is often used instead of a combined simulation and optimisation approach that a-priori integrates the stakeholders goal orientation and risk management into a best-fit design solution. However, recently developed state-of-the-art planning and control methodologies that do already incorporate this integration still lack a stochastic representation and/or an associative multi-objective approach. This paper presents a new project management methodology, called Odycon (Open design and dynamic control), which dissolves the aforementioned shortcomings. To enable this, a generic mathematical statement is first formulated for project planning and dynamic control uniting technical (logistical) capabilities, human goal-oriented behaviour, and stakeholders interests. Then, both Monte-Carlo simulation (MCS) and the Integrative Maximisation of Aggregated Preferences (IMAP) optimisation method are combined, resulting in a best-fit strategic planning and dynamic control methodology. Odycon’s use and added value are demonstrated for two applications: (1) a pure strategic planning of an offshore wind installation project, and (2) a pure dynamic control of a highway infrastructure construction project. Both applications demonstrate Odycon’s advances in concurrent and associative design and decision-making, offering best-fit for common purpose synthesis for different complex project phases.
keywords:
Project management , Project scheduling , Project control , Risk management , Multi-objective optimisation , Monte Carlo simulation , Preference function modelling
journal: arXiv
[…]
[a]organization=Data Science & Engineering, Boskalis,addressline=Rosmolenweg 20, city=Papendrecht, postcode=3356 LK, state=Zuid-Holland, country=The Netherlands
[…]
[b]organization=Corporate R&D, Boskalis,addressline=Rosmolenweg 20, city=Papendrecht, postcode=3356 LK, state=Zuid-Holland, country=The Netherlands
1 Introduction
Successful project delivery is traditionally attempted through project planning and control, answering what is going to be done when and how much does it cost. Planning is referred to the process of deciding how to execute a project, while project control reflects the process of ensuring execution as planned. Establishing a robust and reliable project schedule is fundamental for preliminary cost and time estimates, as well as for effective and efficient project management (Smith et al., 2014). However, with increasing size and degree of uncertainty, initial estimates of costs and timelines tend to be fraught with inherent risks and uncertainty that cannot be completely eliminated. These uncertainties often results in discrepancies between early predictions and stakeholder objectives, complicating the decision-making process and may result in unmet expectations. Initial project planning and execution control is often approached by individual optimisation or simulation methods, resulting in a one-sided point of view and neglecting the multifaceted nature of projects where the technical capabilities, the human goal-oriented behaviour, and the stakeholders desirability are interconnected. It requires a systems-oriented approach to acknowledge these complexities and recognise the need for stakeholder-oriented decision-making to answer what should be accomplished why and how (Reschke and Schelle, 2013).
The role of well-informed human decision-making is crucial in project management, where decision-makers must understand the rationale behind their chosen project planning or control strategies towards risk management, maximum effectiveness, and ultimately stakeholder satisfaction. When evaluating the overall project’s success, it becomes clear that factors beyond just cost and timely completion play a significant role (Van Gunsteren, 2011). Trade-offs between quality, costs, and schedule are a constant reality in construction projects and are typically determined by the involved stakeholders. These trade-offs are reflected by the socio-technical interplay between the stakeholders preference (human domain) and the capability of the project (technical domain), see Van Heukelum et al. (2024). However, the traditional best practice recommends solely adhering to the contract and making a plan without considering the actual project reality, often following the rule to make a plan and stick to it. This practice can be conceived as ’pseudo-reality’, as project success is created by the dynamic adaptation to issues that arise (most often not defined in a contract) and dynamic interaction between stakeholders towards associative solutions. Despite this reality, the traditional managerial focus in construction projects often leans heavily towards costs and schedule and their individual interest, thus neglecting the interplay with stakeholder-oriented concurrent objectives in the process. When stakeholders are willing to go beyond their pure self-interest, joint project success can be enhanced. Stakeholders do have to relinquish their identified individual goals, ultimately enabling a synthesis solution that best fits the common purpose and outperforms single-sided project or even compromise solutions (Glasl, 1998; Scharmer, 2016; Van den Doel and Van Velthoven, 1993; Wolfert, 2023).
Traditional planning & control methods lack to explicitly express the relationship and influence of all participating stakeholder interests towards the decision-making process in project or operations management (Blanchard and Fabrycky, 2011; Del Pico, 2023; Slack et al., 2010). Usually (both in science and in industry), these methods are simulation and analysis oriented (a-posteriori) and do not take an (a-priori) design approach to effectively and efficiently generate a best-fit management solution. However, state-of-the-art methods that are optimisation-based often do not reflect a best-fit for common purpose consideration. These are most often being limited to a single objective optimisation (SOO) approach, relying on monetisation to reflect multiple objectives and neglecting the complexity of stakeholder satisfactory project management. While the benefit of multi-objective optimisation (MOO) in project management has been acknowledged (Guo and Zhang, 2022), there has been no methods developed towards adapting MOO in strategic project planning and dynamic project control integrating probabilistic risk management (Kammouh et al., 2022). Well-informed decision-making is facilitated by removing overconfidence bias through the use of structured and data-driven decision-making processes (Kahneman, 2011). Considering the inclusion of the goal-orientation of all involved stakeholders and their desired expectations, the increasing complexity of construction projects necessitates effective and efficient computer-aided decision support for stakeholder-oriented project management.
To address the lack of consideration for human control and adaptive behaviour in standard probabilistic scheduling methods using Monte-Carlo simulation (MCS) (Del Pico, 2023), research by Kammouh et al. (2022, 2021) has proposed the mitigation controller (MitC), a current state-of-the art project control software. It automates the project manager’s selective nature to only apply the most effective and efficient control measures on-the-run, based on the unique circumstances of each possible project realisation. The MitC, is designed as a SOO control tool, optimising for cost-effectiveness given a project delivery date (target constraint). As a result, there is no understanding of how the most optimal set of control measures impacts other project objectives from relevant decision-makers. While other objectives can be assessed independently, there is no guarantee that an optimal project planning and control strategy is derived that represents a best-fit for common purpose synthesis. Moreover, with increasing project complexity, the impact of individual planning & control variables related to all stakeholder objectives becomes cumbersome to be evaluated in a manual trial-and-error approach.
Associative project planning and control enabling a stakeholder-oriented behaviour can be introduced by utilising the advances developed in the Open Design Systems (Odesys) methodology (Wolfert, 2023), that enables the association of the individual stakeholder interests (desirability) into the technical design optimisation (capability) using integrative maximisation of aggregated preferences (IMAP). This Odesys/IMAP approach proves to be an effective methodology to arrive at best-fit for common purpose design solutions where all stakeholder interest are accounted for and translated to a common preference domain to find an overall group optimum. The introduced IMAP method enables iterative group design and decision-making in a deterministic manner. However, to effectively utilise these advancements for project planning and control, it is crucial to incorporate traditional probabilistic project planning and control methods that can precisely capture and represent the project’s schedule capabilities, such as the Program Evaluation and Review Technique (PERT). Additionally, these tools should employ stochastic simulation methods to reflect the inherent uncertainties and unpredictability encountered both in the planning and execution phases.
Utilising and combining these advances in simulation and optimisation can notably reduce inefficient and misinformed decision-making in project planning and control, particularly with increasing size of projects, where it becomes virtually impossible to account for all possible scenarios. In this context, the focus of this paper is to introduce a methodology that integrates (1) the physical and logistical capability of the project (’network performance domain’), with (2) the human goal-oriented behaviour (’decision-maker objective domain’) and extends it towards (3) the desirability of the project stakeholders (’associative preference domain’) enabling best-fit for common purpose project management. This is required as the current state-of-the-art methodologies for project planning or control exhibit the above-mentioned shortcomings that make them unsuitable for delivering the necessary integrative & associative behaviour to successfully execute the project so that it best meets all stakeholder interests. To achieve such a best-for-project fit, the project management practice PII introduced by Van Gunsteren (2011) and based on Schön and Argyris (1996) can be further improved to a PIII level. This advancement goes beyond the integrative double-loop learning approach, which involves making decisions for the stakeholder or project client. Instead, PIII emphasizes open-loop learning by working closely in association with all project stakeholders to achieve the best possible overall project result. A step from stakeholder oriented, but static ’self’-planning and control, to pure dynamic ’group’-planning and organisation where an integrative maximization of associated preferences is achieved Wolfert (2023).
Based upon the identified development gap and utilisation needs, this paper presents a new integrative stochastic simulation and multi-objective planning & control optimisation methodology, named Odycon (an acronym of open design and dynamic control). Odycon enables best-fit for common purpose generation of both strategic planning and/or dynamic control applications. Odycon is a decision-support methodology that automates the selection of planning and control variables, taking into consideration multiple stakeholder objectives and constraints. Building upon the MitC introduced by Kammouh et al. (2021) and the IMAP optimisation method introduced by Wolfert (2023) and Van Heukelum et al. (2024), this paper aims to advance open and dynamic project planning and control of complex construction projects, fostering multi-objective stakeholder-oriented decision-making. The following advances are presented:
(1) A general framework for an integrated strategic planning and dynamic control problem, which combines probabilistic project schedule simulations with a multi-objective design optimisation. This allows for dynamic adaptation to project-specific needs and stakeholder objectives, enhancing associative decision-making in complex construction projects.
(2) A direct representation between the project feasibility, project management objectives, and stakeholder decisions, achieving a unified integration of network performance, stakeholder objectives, and associative decision-making.
(3) A pure preference-based aggregation for multi-objective decision-making for project management, which translates various objectives and resources into a common preference domain to achieve an optimised group synthesis. This resolves the need for monetisation of non-cost-related objectives and does not go for the lowest cost but for the highest value.
This paper first introduces the general mathematical statement of the multi-objective project planning & control optimisation. Following this, the underlying framework and concept of the stochastic simulation & optimisation methodology are described, with a detailed elaboration on the design of the method, in which the IMAP optimisation is implemented. The methodology is then demonstrated using a pure strategic planning example (no control variables) for an offshore wind installation project, where a service provider and a marine contractor are the key decision-makers. In contrast to the example focusing on pure strategic planning, the paper also demonstrates a pure dynamic control example (no planning variables) for a highway infrastructure project. This example includes associative control by both a contractor and a highway agency, showcasing the application of the methodology for two distinct project management phases. Finally, recommendations for further development are provided, and the added value of the newly developed stochastic simulation and optimisation methodology is concluded.
2 Mathematical formulation of the multi-objective optimisation
To address the mentioned shortcomings of project planning & control, a new stochastic optimisation & simulation methodology is established that enables the intricate representation of all stakeholders interests towards the different project objectives. Reflecting the goal-oriented behaviour of involved stakeholders is achieved by integrating the human preference domain (associated stakeholder interests) and the physical planning performance domain using the Odesys methodology embedded within a stochastic simulation framework (Wolfert, 2023; Van Heukelum et al., 2024). In other words, design-oriented stakeholder behaviour needs to be integrated into probabilistic planning performance. To this end, the integrative multi-objective design optimisation method IMAP, as part of the Odesys methodology, must first be formulated for a multi-objective planning & control optimisation problem. This threefold integration of performance-, objective- and preference functions, reads as follows:
$\begin{gathered}\mathop{Maximise}_{\mathbf{x}}\textrm{ }U=A\left[P_{k,i}\left(O_{i}\left(\mathcal{N}(\mathbf{x},\mathbf{y}),F_{1}(\mathbf{x},\mathbf{y}),F_{2}(\mathbf{x},\mathbf{y}),...,F_{J}(\mathbf{x},\mathbf{y})\right)\right),w^{\prime}_{k,i}\right]\textrm{ for }\\ k=1,2,...,K\\ i=1,2,...,I\end{gathered}$ (1)
Subject to:
$\begin{gathered}g_{p}(O_{i}(\mathcal{N}(\mathbf{x},\mathbf{y}),F_{1,2,...,J}(\mathbf{x},\mathbf{y})),F_{1,2,...,J}(\mathbf{x},\mathbf{y}))\leq 0\textrm{ for }p=1,2,...,P\end{gathered}$ (2)
$\begin{gathered}h_{q}(O_{i}(\mathcal{N}(\mathbf{x},\mathbf{y}),F_{1,2,...,J}(\mathbf{x},\mathbf{y})),F_{1,2,...,J}(\mathbf{x},\mathbf{y}))=0\textrm{ for }q=1,2,...,Q\end{gathered}$ (3)
With:
  • 1.
    $U$: Utility function that needs to be maximized for a best-fit configuration, is addressed using the Genetic Algorithm software tool, preferendus (Van Heukelum et al., 2024).
  • 2.
    $A$: Aggregated preference score determined as part of the IMAP optimisation method (see A).
  • 3.
    $P_{k,i}(O_{i}(\mathcal{N}(\mathbf{x},\mathbf{y}),F_{1,2,...,J}(\mathbf{x},\mathbf{y}))$: Preference functions that describe the preference stakeholder $k$ has towards objective function $i$, which are functions of different performance functions and dependent on planning and/or control variables and physical variables ( $i\leq k$ and $K$ is the maximum number of stakeholders).
  • 4.
    $O_{i}(\mathcal{N}(\mathbf{x},\mathbf{y}),F_{1,2,...,J}(\mathbf{x},\mathbf{y}))$: Objective functions that describe the objective $i$, functions of different performance functions, planning and/or control variables and physical variables.
  • 5.
    $\mathcal{N}(\mathbf{x},\mathbf{y})$: Performance function that describes the project network, depending on planning and/or control variables $\mathbf{x}$ and physical variables $\mathbf{y}$. Here $\mathcal{N}$ is a logical representation of a physical project planning network, composed of nodes (activities) and edges (logical links).
  • 6.
    $F_{1,2,...,J}(\mathbf{x},\mathbf{y})$: Performance functions that describe the object behaviour, depending on the planning and/or control variables $\mathbf{x}$ and (constant) physical variables $\mathbf{y}$.
  • 7.
    $\mathbf{x}$: A vector containing the available planning & control variables (i.e. endogenous design variables) $x_{1},x_{2},...,x_{N}$. These variables are bounded such that $lb_{n}\leq x_{n}\leq ub_{n}$, where $lb_{n}$ is the lower bound, $ub_{n}$ is the upper bound.
  • 8.
    $\mathbf{y}$: A vector containing the physical variables (i.e. exogenous ’non-design’ variables) $y_{1},y_{2},...,y_{M}$.
  • 9.
    $g_{p}(O_{i}(\mathcal{N}(\mathbf{x},\mathbf{y}),F_{1,2,...,J}(\mathbf{x},\mathbf{y})),F_{1,2,...,J}(\mathbf{x},\mathbf{y}))$: Inequality constraint functions, which can be either objective function and/or performance function constraints.
  • 10.
    $h_{q}(O_{i}(\mathcal{N}(\mathbf{x},\mathbf{y}),F_{1,2,...,J}(\mathbf{x},\mathbf{y})),F_{1,2,...,J}(\mathbf{x},\mathbf{y}))$: Equality constraint functions, which can be either objective function and/or performance function constraints.
  • 11.
    $w^{\prime}_{k,i}$: Weights for each of the preference functions (sum of the number of preference functions equals $K*I$). The global weights for the relative importance of stakeholders is defined as $w_{k}$. The local weight stakeholder $k$ gives to objective $i$ is defined as $w_{k,i}$. The following formula holds: $w^{\prime}_{k,i}=w_{k}\cdot w_{k,i}$, given that $\sum w^{\prime}_{k,i}=\sum w_{k,i}=\sum w_{k}=1$. Note that in case of equivalent stakeholder decision-making $w_{k}=1/K$.
To further elaborate on this mathematical formulation, several important remarks are made which are discussed below.
Remark 1: Preference aggregation
The preference aggregation is performed using the principles of the PFM theory (Barzilai, 2022). It is an integral part of the integrative preference maximisation method (IMAP), see Wolfert (2023). To arrive at an aggregated preference for individual objectives, the IMAP method employs the mathematical operator, A, which determines the aggregated preference score using an algorithm called the A-fine Aggregator (see A).
Remark 2: Preference functions
The preference functions describe the relation of a stakeholders satisfaction towards a certain objective on a defined scale. The elicitation of the preference functions and associated individual weights is essential to reflect the relationship of the stakeholders interest towards a certain objective. As an initial estimation for both the preference functions input and local and global the weights the Conjoint Analysis (CA) method can be used. For more information regarding the process of preference elicitation, the reader is referred to Arkesteijn et al. (2017). As mentioned in Van Heukelum et al. (2024) objective $O_{i}$ can be associated with multiple stakeholder preference functions $P_{k,i}$ (as $k\geq i$). Therefore, it is not required that a stakeholder expresses a preference for all objectives.
Remark 3: Network
Optimisation requires representing the network planning as a Directed Acyclic Graph (DAG), which is a structure composed of nodes (project activities) and directed edges (activity interdependencies: i.e, logical links) that do not form any cycles, ensuring all paths from the initial node to any other node are open-ended and loop-free.
Remark 4: Planning & control variables
In the context of a pure project planning optimisation (no control measures), the variables $x_{n}$ are defined as stated above, allowing to take any form (e.g. integer, float, binary, etc.). However, for a pure project control optimisation (no planning variables), the variables $\mathbf{x}$ are redefined as $\mathbf{x}=\mathbf{a}\mathbf{c_{i}}$, that is the product of the allocation of a control variables vector $\mathbf{a}$ (which takes binary value $a_{n}\in\{0,1\}$, where 0 reflects not allocated, or 1 reflects allocated) and the control impact vector $\mathbf{c_{i}}$ towards objective $i$, containing the impact of control variables $c_{1},c_{2},...,c_{N}$ directly related towards control variables $x_{1},x_{2},...,x_{N}$.
3 Odycon’s stochastic simulation & optimisation methodology
This section describes the conceptual functioning of the Odycon methodology, with its underlying stochastic simulation & optimisation framework combining probabilistic MCS and IMAP multi-objective optimisation to enable associative project control for best-fit for common purpose project execution. The essence is to account for the inherent uncertainties of a project combined with the goal-oriented objectives of all stakeholders. The following two sections elaborate on the integrated MCS and the utilised IMAP optimisation method.
3.1 Probabilistic MCS approach
To accurately reflect uncertainties and unpredictability encountered during project execution, the core of the Odycon methodology is build upon a probabilistic MCS frequently used for modelling uncertainties and capturing the stochastic behaviour of variables in project scheduling (Del Pico, 2023). The duration uncertainty of activities, impact of risks, impact of other external influences, or control measures is quantified by the Beta-PERT distribution. This distribution is widely recognised for modelling uncertainties in project scheduling. Accordingly, the probability density functions for the varying parameters are denoted as $f(z_{i};a_{i},m_{i},b_{i})$, where $a_{i}$, $m_{i}$, and $b_{i}$ represent the optimistic, most likely, and pessimistic time estimates, respectively.
Within an MCS iteration, random values $z_{i}$ will be sampled according to the given respective probability distributions. The occurrence of a risk is modelled by a binary variable $X$, where a risk occurs $X=1$ with probability $p_{e}$, and a risk does not occur $X=0$ with probability $1-p_{e}$. In deterministic project scheduling (using the critical path method), time estimates are fixed, resulting in one static project network realisation. In contrast, with MCS used in probabilistic project planning (PERT), the structure of the project network $\mathcal{N}$ changes with each iteration based on random sampling, creating various realisations of project networks. This variability reflects the inherent uncertainty and stochastic nature modelled in the MCS, allowing for a comprehensive analysis of frequency distribution of optimal project planning and control strategies. This approach also aids in identifying the most probable project paths, critical activities, thereby enhancing decision-making and risk management strategies in project planning and execution.
The framework enables decision-making through the MCS given the steps provided in Figure 1. (1) First, the Project network data is read, and the network $\mathcal{N}(\mathbf{x},\mathbf{y})$ is established given the project activities, risk event and any other relevant data. Note, to that extent any network planning or simulation method can be integrated, that is able to reflect the necessary logical links in the network. (2) Next, the MCS is initiated by a loop of n runs. (3) Step 3 is to sample random duration’s for the project activities, risk and other external influences (like weather). To establish the optimisation, first the type of variables must be defined. If control variables are not considered, the planning variables are defined and the optimisation initiated (continues with step 6). (4) If control variables are included, the impact $\mathbf{c_{i}}$ towards objective $i$ needs is calculated given the duration capacity of the control measure. (5) Next, the network is compiled and the delay with respect to the target duration evaluated. In case of a delay, the optimisation is initiated. (6) The optimisation is run by defining the objective functions, the preference functions and their related weights, in accordance with the prior established mathematical formulation for the optimisation (see section 2). Once defined, the optimisation is performed utilising the IMAP method. To arrive at the highest aggregated preference score $A$, the values of the objectives functions are calculated and with that the corresponding value of the preference functions. Once established, the aggregated preference score can be determined according to the principles of references aggregation with preference maximisation (more information see the section on the optimisation below). (7) After every iteration of the MCS the result is stored. (8) Once the MCS reaches the $n^{th}$ iteration, the results of the individual optimisations can be presented with the criticality index (how often does a certain variable or a combination of variables occur). Given the collected data, detailed statistical analysis are enabled.
3.2 Integrative multi-objective optimisation approach
To integrate the human goal-oriented behaviour in project management, and associative decision-making the multi-objective optimisation formulation (defined in section 2) is integrated into the MCS as part of the IMAP optimisation method. The optimisation is executed within every iteration of the probabilistic MCS, aiming to select the most effective set of project planning variables and or control measures given the specific realisation of the network within that iteration.
The essence of the IMAP method lies in combining preferences aggregation (reflecting associative decision-making) with preference maximisation (reflecting human goal-oriented behaviour). To arrive at the highest aggregated preferences the weighted (or relative) scores of the individual preference functions ($P_{k,i}$) are aggregated in their affine space according to the basic principles of PFM theory. The final preference score aggregation is performed using the weighted least squares method. The preference score aggregation as part of the IMAP optimisation is described in A. To facilitate the application of the IMAP method in practice, the Python-based software tool, preferendus, is incorporated into the framework (Van Heukelum et al., 2024).
With utilising the IMAP optimisation method the shortcoming of classical MOO methods that consider the Pareto front as a valid outcome is addressed. MOO methods using the Pareto front result in a large set of multiple possible solution points, with the need to evaluate these solution points a posteriori. Moreover, given the complexity of a project the set of planning & control variables and objectives can become very large, consider a for example a project with 4 stakeholders and 40 control variables. In such a case, project managers and stakeholder are interested in a single best design solution. We want to avoid cumbersome a posteriori evaluations of possible suboptimal solutions, resulting from a trial-and-error approach. Classical Pareto front based methods are in this planning & control approach case inappropriate (see Van Heukelum et al. (2024)). IMAP instead results in an integrative a priori optimisation with a single best-fit result from a multidimensional solution space.
Figure 1: Odycon concept diagram.
4 Demonstrative planning & control applications
The introduced Odycon methodology is demonstrated in this section upon two real-life multi-stakeholder construction projects. To demonstrate the Odycon advances towards concurrent and associative design and decision-making, two examples are presented: (1) a pure strategic planning application (no control variables) for an offshore wind installation project, and (2) a pure dynamic control application (no planning variables) of a highway infrastructure construction project.
To legitimise the choice of these two examples, the following is considered. In general, two types of project management approaches are defined. For projects governed by strict procedural continuity with limited flexibility to deviate from planned tasks (the so-called Toothpaste model), management tends to prioritise strategic planning in advance. In contrast, for projects that allow considerable flexibility and adjustments during implementation (the so-called Viking model), management strategies allow primarily for dynamic control on-the-run (Reschke and Schelle, 2013).
Unlike inland highway projects where extra assets are often readily available and can be mobilised during project execution, offshore projects require detailed front-end planning due to the limited possibility of on-the-run control. For this reason, a pure planning scenario for this ’Toothpaste’ project is considered where the focus is to select the most impactful strategies up-front, reducing the potential risk during execution. To show the effect of dynamic control on-the-run, a pure control scenario for a ’Viking’ project is chosen, which is demonstrated on the basis of a transport infrastructure project. These demonstrative applications illustrate how the introduced Odycon methodology can support optimal decision-making for execution and mitigation strategies. The complexity of these projects demands computer aided design and decision-making, where the participating stakeholders together decide to arrive at best-fit for common project purpose.
To this end, two Python based models were developed: one for a purely strategic planning case (SYPL, an acronym of strategic planning) and another for a purely dynamic control case (MICO, an acronym of mitigation control). The source code, project data, and results of these demonstrative applications are available on GitHub, as detailed in the data availability statement.
Final note, both application examples are based upon previous publications, all of which one of the authors of this paper was involved in, see Van Heukelum et al. (2024) and/or Kammouh et al. (2022). To read these examples here independently, the project descriptions are partly repeated.
4.1 Strategic planning application
Technical context: Offshore floating wind (OFW) technologies are considered a solution for wind energy production in deep waters and regions with stronger wind velocities (Spring, 2020). Contrary to the conventional method of employing fixed monopiles for bottom-founded installations, these turbines are mounted on platforms that are anchored to the seabed using mooring lines. The OFW site consists of 36 floating wind turbines (FWT) and 108 suction anchors (i.e., 3 anchors per FWT). OFW projects are executed by a set of highly specialised vessels, where the installation of the individual foundations or floaters is often performed in an iterative or cyclic process (serial installation). The cyclic installation process is often constrained by a singular, inflexible execution path, leaving little room for deviation in the face of delays from adverse weather, technical issues, or operational risks. This rigidity necessitates direct resolution of problems, pausing the project if needed, rather than employing alternative strategies.
Social context: This application includes the following concurrent stakeholder objectives both from an energy service provider (Stakeholder 1) and a marine contractor (Stakeholder 2): (1) minimising project duration, (2) reducing installation costs, (3) optimising fleet utilisation, and (4) lowering CO2 emissions. While the energy service provider prioritises fast project completion to expedite revenue generation and aims to decrease CO2 emissions for better environmental and societal acceptance, the marine contractor focuses on cutting costs to enhance competitiveness and seeks to improve fleet utilisation for greater operational efficiency.
We will now first describe the integrative planning optimisation problem by working through the mathematical statement (see section 2), resulting in performance-, objective-, and preference functions.
4.1.1 Performance functions
Planning variables
The installation of OFW turbines requires specialised types of installation vessel. The performance of the planning sequence is dependent on the utilisation of these vessels during construction. The type of vessels and their properties used in the project are defined in Table 1. The boundaries are the (integer) number of possible vessel to be utilised. The utilisation refers to the probability of a vessel being better utilised in a different project of the contractor.
Table 1: Strategic planning application: Planning variables (OCV refers to offshore construction vessel).
Variables Description Boundaries Anchor deck space Day rate $R_{i}$ utilisation prob. $p_{i}$ CO2 emissions $E_{i}$
$F1=x1$ small OCV $0\leq x1\leq 3$ 8 $47k\text{\texteuro}$ 0.7 30 $30\text{\ }t/day$
$F2=x2$ large OCV $0\leq x1\leq 2$ 12 $55k\text{\texteuro}$ 0.8 40 $40\text{\ }t/day$
$F3=x3$ barge $0\leq x1\leq 2$ 16 $35k\text{\texteuro}$ 0.5 35 $35\text{\ }t/day$
Project activities
The installation process involves strategically placing and securing mooring anchors using a combination of vessels, adjusted for efficiency based on available deck space and potential risk event delays. Two distinct activities can be defined: (1) installing the available anchors on deck of the vessels; and (2) bunkering, which occurs after the installation sequence once all available anchors have been installed. Bunkering refers to the transfer of additional anchors to the vessels when the condition that the number of anchors left to install is zero or less is met.
The duration of these two activities are defined in Table 2 and represented by the pessimistic $a_{i}$, most-likely $m_{i}$, and optimistic $b_{i}$ time estimates. These values are usually obtained either from experience or past data of similar projects and make up the three-point estimates of the activities duration’s, which are used to define the Beta-PERT distributions (see section 3).
Table 2: Strategic planning application: Project activities
Activity description Activity Duration (days)
Optimistic $a_{i}$ Most-likely $m_{i}$ Pessimistic $c_{i}$
Installation 0.80 1.00 1.20
Bunkering small OCV 1.20 1.50 1.80
Bunkering large OCV 1.60 2.00 1.40
Bunkering barge 2.00 2.50 3.00
Risk events
The offshore installation sequence increases in complexity due to uncertainties in the execution. External weather impacting the workability window of vessels, and project specific constraint often disrupt the execution, making completion according to initial forecast difficult to achieve. In probabilistic planning, risk events are included by considering their occurrence probability. Table 3 defines the risk events identified with their corresponding three-point duration estimates. The probability density $f(z_{i};a_{i},m_{i},b_{i})$ can be built using the three-point estimate. For every risk event, an occurrence probability $p_{r}$ is defined.
Table 3: Strategic planning application: Risk events.
Risk Event Risk duration [$days$] Affected Activity Probability $p_{r}$
$a_{i}$ $m_{i}$ $b_{i}$
Weather Delay 0.50 1.00 1.50 Installation 0.20
Supply Chain Issue 1.00 1.50 2.00 Bunkering 0.10
Technical Failure 0.50 1.00 1.50 Installation 0.15
Logistical Constraints 1.00 1.50 3.00 Bunkering 0.20
Marine Traffic 0.50 0.75 2.00 Installation 0.05
Environmental Restrictions 2.00 3.50 5.00 Bunkering 0.10
Lack of Skilled Personnel 1.00 2.00 3.00 Installation 0.08
Equipment Shortage 1.00 2.00 4.00 Installation 0.12
The installation sequence (network) is represented and established using a discrete event simulation (DES), which depends on the planning variables, activities and risks defined above. The probabilistic network with the underlying logical links is expressed as follows:
$\mathcal{N}(x_{1},x_{2},x_{3},\mathbf{y})$ (4)
where the physical variables $\mathbf{y}$ contain the information on the respective vessel deck space. To ensure that the project is executed by at least one vessel, the sum of all vessels on the project must be greater than one. The technical feasibility is guaranteed by a constraint ensuring the anchor’s resistance meets or exceeds the applied forces defined by further performance functions ($F_{4}$ ,…, $F_{7}$). For a thorough exposition of these formulations and their underlying principles, the reader is referred to Wolfert (2023) or Van Heukelum et al. (2024).
4.1.2 Objective functions
The optimisation framework considers the following four objectives that form the link between the network performance function and the preference functions. For a thorough exposition of these formulations and their underlying principles, the reader is referred to Wolfert (2023) or Van Heukelum et al. (2024). The following objective functions are rewritten based upon to suit the needs of the Odycon framework.
  1. 1.
    Project duration: the project duration $\Delta$ extracted from the network performance function expressed as:
    $O_{1}=O_{PD}=\Delta(\mathcal{N}(x_{1},x_{2},x_{3},\mathbf{y}))$ (5)
  2. 2.
    Installation cost: the project cost depends on the day rates of the vessels $R_{i}$ multiplied with their active duration on the project $t_{i}$ given the network performance function, and the cost of the installed anchors, expressed as:
    $O_{2}=O_{C}=(815M_{a}+40,000)n_{a}+\sum_{i=1}^{3}x_{i}R_{i}t_{i}(\mathcal{N}(x_{1},x_{2},x_{3},\mathbf{y}))$ (6)
  3. 3.
    Fleet utilisation: utilisation is represented by the probability of a vessel being better utilised in another project and is defined as:
    $O_{3}=O_{F}=\prod\limits_{i=1}^{3}p_{i}^{x_{i}}$ (7)
  4. 4.
    Sustainability (CO2 emissions): emissions are defined as the daily vessel emissions $E_{i}$ multiplied with their active duration on the project $t_{i}$ given the network performance function:
    $O_{4}=O_{S}=\sum_{i=1}^{3}x_{i}E_{i}t_{i}(\mathcal{N}(x_{1},x_{2},x_{3},\mathbf{y}))$ (8)
4.1.3 Preference functions
The preference functions for this demonstrative application are defined based on the input from industry project management professionals. The maximum preference ($P_{1,PD}=100$) towards the objective $O_{PD}$ is defined as the target duration $T_{tar}=90$ days. A Beta-PERT distribution is employed to model the preference towards project duration. This unique preference modelling reflects that, in principle, the project manager in reality is somewhat interested in an earlier project delivery than the exact target duration, but certainly not much later than this duration with a bit of slack. If there is a greater interest to deliver earlier, then the area under the preference function should be increased so that a more asymmetric function is resulting. This approach enables a continuous evaluation of preferences towards the project duration, facilitating more informed and balanced management decisions amidst uncertainty. The other three preference functions for cost, fleet utilisation and CO2 emissions are respectively convex or concave functions with the max preference value ($P_{1,C}=P_{2,F}=P_{2,S}=100$) for the lowest respective objective values (min $O_{i}$) and with the lowest preference value ($P_{1,C}=P_{2,F}=P_{2,S}=0$) for the highest possible objective values (max $O_{i}$) respectively. The resulting preference functions, describing the relations between the different preferences and the objectives, are depicted in Figure 2.
4.1.4 Results
To retrieve outcomes, we will first have to estimate the weights to generate IMAP solutions as part of the simulation & optimisation framework. The global weights $w_{k}$ are split equally between the energy provider ($w_{1}=0.5$) and the marine contractor ($w_{2}=0.5$). To reflect the objectives of their interest, their individual local weights are respectively: $w_{1,PD}=0.60$ for project duration, $w_{1,S}=0.40$ for CO2 emissions, $w_{2,C}=0.70$ for the installation costs, and $w_{2,F}=0.30$ for fleet utilisation. Table 4 defines the weights of each of the preference functions.
Table 4: Weights for each of the preference functions, according to $w^{\prime}_{k,i}=w_{k}\cdot w_{k,i}$. Note, stakeholders can also reflect interest to all four objectives while the global weight distribution is constant.
Stakeholder $k$ $w^{\prime}_{k,PD}$ $w^{\prime}_{k,S}$ $w^{\prime}_{k,C}$ $w^{\prime}_{k,PD}$ $w_{k}$
Energy provider 0.30 0.20 0.00 0.00 0.50
Marine contractor 0.00 0.00 0.35 0.15 0.50
Total 0.30 0.20 0.35 0.15 1.00
The final outcomes of the Odycon framework, are displayed in the form of frequency distribution diagrams referred to as criticality index. Within every MCS iteration, a new set of vessel combinations (planning variables) is derived through the IMAP optimisation. Given the unique realisation of the network, the vessel combinations change from one iteration to another. To identify the most critical set of planning variables, Figure 3 gives insight into the frequency of number of vessels per vessel type and the frequency distribution of optimal planning variables during the optimisation for both evaluation cases. The outcome reflects the percentage of MCS iterations of which a specific set of vessel combinations is derived. Table 5 compares the numerical results for the different evaluations given defined target-percentiles. The outcomes of the different percentiles per optimisation method are plotted in the different preference functions showing the respective preference given the percentile (see Figure 2). Note, as the MCS is stochastic in nature, it is not possible to provide a single (deterministic) outcome.
Figure 2: The four preference functions ($P_{1,PD},P_{1,C},P_{2,F},P_{2,S}$) with respect to objectives ($O_{PD},O_{C},O_{F},O_{S}$), including the preference score of the different optimisations and percentiles. The numerical results can be found in Table 5.
Table 5: Evaluation of cases with 2000 MCS iterations ($O_{PD}$ in days, $O_{C}$ in $10^{7}$, $O_{F}$ in number of vessels [-], $O_{S}$ in tonnes). Where the MOO case considers $w^{\prime}_{1,PD}=0.30,w^{\prime}_{1,S}=0.20,w^{\prime}_{2,C}=0.35,w^{\prime}_{2,F}=0.15$.
Case P50 P80 P90
$O_{PD}$ $O_{C}$ $O_{F}$ $O_{S}$ $O_{PD}$ $O_{C}$ $O_{F}$ $O_{S}$ $O_{PD}$ $O_{C}$ $O_{F}$ $O_{S}$
SOO $O_{C}$ 118 1.08 0.07 8152 130 1.13 0.35 9000 135 1.15 0.50 9380
MOO 88 1.14 0.04 7300 93 1.17 0.35 8700 95 1.19 0.35 9000
abs. diff. 30 -0.06 0.03 852 37 -0.04 0.00 300 40 -0.04 0.15 380
rel. diff. 25.4 -5.6 42.9 10.5 28.5 -3.5 0.0 3.3 29.6 -3.5 30.0 4.0
For comparative evaluation purposes, the MOO with the earlier defined weights and preferences is compared to a traditional SOO approach towards minimal cost (SOO $O_{C}$). The following conclusions can be made from the outcome:
  1. 1.
    The SOO $O_{C}$ approach results in slightly lower costs for all percentiles (avg. 4% lower) compared to the MOO approach by utilising no small OCV in 80 % of the simulations, suggesting the cost-efficiency of the small OCV to be low (see Table 5). It becomes evident that more vessels lead to higher costs.
  2. 2.
    Comparing the SOO $O_{C}$ with the MOO approach, it becomes evident that considering project cost alone will not result in configurations that lead to overall favourable outcomes. While the project costs differ very little between the approaches, the MOO outperforms the SOO $O_{C}$ for all other objectives and percentiles leading to lower project duration’s (avg. 28% lower), lower probability of the fleet being better utilised elsewhere (avg. 24% lower), and lower emission rates (avg. 6% lower). This can be achieved by introducing small OCVs. This is explained by the reflection of interests towards the sustainability and fleet utilisation in the MOO. This comparison highlights that focusing solely on cost (single objective approach) does not accurately capture the complexities of stakeholder-oriented project planning. In contrast, utilising the MOO approach offers a more balanced perspective by taking into account all stakeholder interests and achieves a higher aggregated preference compare to the SOO $O_{C}$ approach, including the inherent uncertainties and stochastic nature of the project leading to overall favourable project realisations.
  3. 3.
    Figure 2 illustrates that as the percentile increases, the preference for both cases decreases concerning their respective objectives. This trend indicates that decision-makers encounter lower preference outcomes when aiming for higher objective probabilities. The target probability can thus not only be reflected by the unit of the individual objectives but also directly in the common preference domain. By examining the target-preference values, decision-makers can identify the probability of achieving a particular outcome.
  4. 4.
    The criticality index (see Figure 3) aids decision-making by clarifying the utilisation of specific types of vessels and their combinations, thereby identifying the most effective vessel configurations given the underlying uncertainties and stochastic factors. For instance, in both optimisation scenarios, large OCVs are not utilised in 80% of the simulations. These observations provide project managers with actionable insights, enabling them to allocate their planning and tendering budgets more effectively towards solutions that are both efficient and feasible.
Figure 3: Criticality index of number of small OCV’s (top left), large OCV’s (top right), barges (bottom left) and vessel combinations (bottom right) for both optimisation cases. Note, only combinations that stored during the simulation & optimization are considered in this figure.
4.2 Dynamic control application
Technical context: The Schiphol-Amsterdam-Almere (SAA) Project is the largest road construction project in the Netherlands in the period from 2012 to 2024. It realises an expansion of the capacity of the Dutch national highways A6, A1, A10-east and A9 and thus aims to improve the accessibility of the northern part of the Randstad (Schiphol, Amsterdam, Almere). Project SAA includes 63 kilometres of motorway, the construction of two tunnels, two large bridges and an aqueduct; the modification of five interchanges, and about 100 engineering structures. As part of this project, the SAAone consortium (Boskalis, Volker-Wessels, Hochtief and DIF), together with the Dutch highways agency Rijkswaterstaat, provided the design, construction, maintenance and operation. The major construction project took more than four years and consisted of a massive schedule that could be regularly adjusted on the run.
Social context: This application includes the following concurrent stakeholder objectives both from SAAone and RWS: (1) meeting target completion, (2) managing project control costs, and (3) minimising the traffic nuisance. In order to manage the project in the best possible way, the costs of the extra control measures to be allocated, which are possibly not part of SAAone’s risk budget, are at odds with the opening of the new highway on time (interest RWS). Even more, steering solely on the completion date from RWS’s project delivery department perspective may also be of opposite interest to minimising traffic nuisance during construction, which RWS’s service operations department is responsible for. That is what makes this example a dynamic multi-objective control and tripartite stakeholder problem.
We will now first describe the dynamic control optimisation problem by working through the mathematical statement (see section 2), resulting in performance-, objective-, and preference functions.
4.2.1 Performance functions
Project activities
The project consists of a number of distinct activities, as detailed in Table 8. For each activity, the durations are presented by an optimal $a_{i}$, most likely $m_{i}$, and pessimistic $b_{i}$ time estimate. These values contribute to the three-point estimates for the activity duration’s, forming the basis for the Beta-PERT distributions. Additionally, the sequence of activities is defined by specifying predecessors for each activity, thereby integrating the dependencies among activities into the network model.
Next to the individual uncertainty of activity duration’s a set of shared uncertainties is considered to model activity duration correlations arising from common factors (e.g., weather, labour skills). The implementation of stochastic activity correlation was implemented in accordance with the methodology and underlining example presented in Kammouh et al. (2022). For further information regarding the evaluation of networks under correlated uncertainty, see Wang and Demsetz (2000). Table 9 defines the uncertainty factors, their three-point estimates and relationships with the project activities.
Risk events
The project is affected by many possible risk events that ad an additional source of uncertainty that can negatively affect the desired project outcome (see Table 10). The impact of these risks is defined by their risk duration (three-point duration estimates) in the same manner as done for the activities and control measures. Every risk is characterised by a given probability of occurrence ($p_{e}$) and the relations between the risk events and activities.
Control measures
Based on the project database, a number of possible control measures $x_{n}$ are identified (see Table 11), that can increase the probability of finishing the project at the target completion time. These control measures, identified by the project manager, take into account limitations on material and human resources. These control measures are characterised by their impact towards the three identified objectives: (1) the number of days reducing the duration of the affected activities, (2) the cost of a measure, and (3) the impact on traffic nuisance, each given by a three-point estimate (minimum, most-likely, and maximum). The correlation factor ($\eta$) is defined for accurately determining mitigation effects in construction delays by accounting for the nonlinear relationship between mitigated duration and effects, varying from direct proportional effects for additional personnel or equipment to non-proportional one-off expenses.
The SAAone construction sequence (see Figure 7), is established using a logical network model and build upon the planning project activities and risks, depending on the control measures. The probabilistic network with the underlying logical links is expressed as follows:
$\mathcal{N}(a_{n}*c_{i,n},\mathbf{y})$ (9)
As noted in remark 4 of section 2 the methodology defines the implementation of a control measure to be represented by allocation of a measure $a_{n}$ multiplied with its impact $c_{i,n}$ towards objective $i$ as following: $F_{n}=x_{n}=a_{n}*c_{i,n}$ (see the set of control measures in Table 11).
4.2.2 Objective functions
The optimisation framework considers the following four objectives that form the link between the network performance function and the preference functions.
Objective project duration
The project duration $\Delta$ is extracted from the network performance function and can be expressed as follows:
$O_{1}=O_{PD}=\Delta(\mathcal{N}(a_{n}*c_{1,n},\mathbf{y}))$ (10)
where $O_{PD}$ is expressed in days.
Objective project control cost
The project control cost is defined as the sum of the cost impact of active control measures. As the demonstrative case is defined according to a Design, Build, Finance and Maintain-contract (DBFM) contract, the objective function is extended by a penalty and reward scheme (contractual project completion performance scheme), it reads as:
$O_{2}=O_{C}=\sum_{n=1}^{N}\left(a_{n}\cdot c_{2,n}\right)+\Delta_{1}\cdot P_{c}-\Delta_{2}\cdot R_{c}$ (11)
Subject to:
$\Delta_{1}=\begin{cases}\Delta-T_{\text{tar}},&\text{if }\Delta>T_{\text{tar}}\\ 0,&\text{otherwise}\end{cases}$ (12)
$\Delta_{2}=\begin{cases}T_{\text{tar}}-\Delta,&\text{if }\Delta<T_{\text{tar}}\\ 0,&\text{otherwise}\end{cases}$ (13)
where $O_{C}$ is expressed in €, $\Delta_{1}$ is the project delay after implementing the control measures, $\Delta_{2}$ is the duration reduction beyond target duration, $T_{tar}$ is the target (i.e., desired/planned) project duration specified by the involved stakeholders, $P_{c}$ is the daily penalty, and $R_{c}$ is the daily reward. It holds that a penalty and reward cannot occur simultaneously, thus $\Delta_{1}*\Delta_{2}=0$. Note, the project duration after implementation of control measures $\Delta$ is extracted from the network performance function $\mathcal{N}(a_{n}*c_{2,n},\mathbf{y})$, see Equation 9.
Objective traffic nuisance
During construction, the experience of a road user (considered as traffic nuisance) is heavily impacted. The impact on road user experience during construction is becoming an increasingly important aspect within infrastructure project execution. Currently, most service operations departments measures the effect on car traffic in terms of Lost Vehicle Hours (LVHs), which is then converted into a monetary value and considered in a contractual performance scheme. However, this approach overlooks the actual impact on the user which can not be monetised, respectively maximising their experience during construction becomes essential for improving service operations during construction. Similar to the penalty and reward scheme of the objective cost, a scheme is implemented to account for nuisance increase with the factor $P_{e}$ in case of a delay and decrease in traffic nuisance with the factor $I_{e}$ in case of early completion. $\Delta_{1}$ and $\Delta_{2}$ are considered according to Equation 12 and Equation 13. Again, it holds that a penalty and reward cannot occur simultaneously, thus $\Delta_{1}*\Delta_{2}=0$. The objective traffic nuisance reads as follows:
$O_{3}=O_{N}=\left(\frac{1}{\sum_{n=1}^{N}\left(c_{3,n}\right)}\sum_{n=1}^{N}\left(a_{n}\cdot c_{3,n}\right)\right)*S+\Delta_{1}*P_{n}-\Delta_{2}*R_{n}$ (14)
where $S$ is the scaling factor of 10. Since there is no standard method on how to represent the traffic nuisance, a scale from 0 (baseline nuisance during construction) to 10 (worst possible traffic nuisance during construction) is introduced.
4.2.3 Preference functions
The preference functions for this project control application were developed with one of the co-authors who was involved in the project during the execution phase. In accordance with the project RWS delivery department and the SAAone consortium, the preference curve for project duration (Time) is defined by three duration estimates using a Beta-PERT distribution. The maximum preference ($P_{1,PD}=P_{2,PD}=100$) towards objective $O_{PD}$ is set at the contract target duration, $T_{tar}=1466$ days. The minimum preference ($P_{1,PD}=P_{2,PD}=0$) is determined by two points: the target duration plus the maximum allowable delay, and a minimum duration of 966 days. Note, the preference towards the project delivery time is expressed by the joint contract of RWS and SAAone, therefore the one preference curve is defined towards the objective target duration. Similar to the prior example, the preference curve is established using a Beta-PERT distribution to accurately reflect the project manager’s preferences for the project duration. This unique preference modelling reflects that, in principle, the project manager in reality is somewhat interested in an earlier project delivery than the exact target duration (blue vertical line), but certainly not much later than this duration with a bit of slack (red vertical line). If the project manager has a ’fixed’ project delivery with no slack, the red and blue line should coincide, essentially reflecting no preference for any delay. If there is an increased interest to deliver earlier, then either the target duration can be changed or the area under preference function should be increased to create an extremely asymmetric function. Since the delay of the project changes with every iteration of the MCS, the preference curve is adapted accordingly for every realisation of the project network. The preferences for lowest project control cost and lowest nuisance are considered to have the highest value ($P=100$). Conversely, the lowest preference value ($P=0$) is considered for the total sum of all control measures’ cost and for the worst possible level of traffic nuisance. For an initial estimate, both preference functions (cost and nuisance) are linear between these lowest and highest preference values. The three resulting preference functions, which describe the relations between different values for $P_{1..3,1..3}$ with respect to objectives $O_{1..3}$, are shown in Figure 4.
4.2.4 Results
To retrieve Odycon outcomes, we will first have to estimate the weights to generate IMAP solutions as part of the MCS results. Within a concurrent discussion between the two (global) stakeholders RWS ($w_{1}=0.5$) and the SAAone consortium ($w_{2}=0.5$) the three individual weights are discussed to reflect their shared objectives. RWS internal interests are represented by the project delivery department with $w_{1,PD}=0.70$ for the project duration and the service operations departments with $w_{1,N}=0.30$ for traffic nuisance. SAAone expressed their individual (local) weights to be $w_{2,PD}=0.70$ for project duration and $w_{2,C}=0.30$ for the control costs. Table 6 reflects the resulting weights of each of the preference functions. As mentioned before, the preference towards the project delivery time is expressed by the joint contract of RWS and SAAone, therefore the one preference curve is defined towards the project duration.
Table 6: Weights for each of the preference functions, according to $w^{\prime}_{k,i}=w_{k}\cdot w_{k,i}$. Note, stakeholders can also reflect interest to all objectives while the global weight distribution is constant.
Stakeholder $k$ $w^{\prime}_{k,PD}$ $w^{\prime}_{k,C}$ $w^{\prime}_{k,N}$ $w_{k}$
RWS 0.35 0.00 0.15 0.50
SAAone consortium 0.35 0.15 0.00 0.50
Total 0.70 0.15 0.15 1.00
For comparative evaluation purposes, three scenarios are considered. First, there is a SOO focused on minimising costs (SOO $O_{C}$). Second, another SOO targets minimising nuisance (SOO $O_{N}$). Finally, a MOO is conducted using a defined set of weights (see Table 6) to reflect all stakeholder’s interests. To ensure convergence towards the project target date $T_{tar}$ a penalty of $P_{c}=10k\text{\texteuro}$ and $P_{n}=0.1$ is defined for SOO costs and SOO nuisance case respectively. Without these ”constraints” the SOO would result in no used control measures due to the missing incentive to utilise them. This is however not needed for the MOO, as it solved the aforementioned need for direct time-cost trade-off techniques.
The final outcomes of the Odycon framework (integration of IMAP and MCS) are displayed in the form of frequency distribution diagrams referred to as criticality index. Similar to the strategic planning application, the criticality index of control measures is derived. Within every MCS iteration a new set of control measures is derived. To identify the most critical set of control variables Figure 5 gives insight into the frequency of control measures and Figure 6 on the frequency of combined occurrence of control measures. Table 7 compares the numerical results for the different evaluation cases given defined target-percentiles. The results of the different percentiles and cases are respectively displayed in the different preference functions showing the respective preference given the percentile (see Figure 4).
Figure 4: Preference functions ($P_{1,PD},P_{2,C},P_{1,N}$) with respect to objectives ($O_{PD},O_{C},O_{N}$), including the preference score of the different optimisation scenarios and percentiles. The numerical results can be found in Table 7. The blue line represents the target duration $T_{tar}=1466$ days and the red line an exemplary maximum delay of 100 days. Note that, as previously mentioned, since the project delay changes with each iteration of the MCS, the preference curve is adapted accordingly for each realisation of the project network. However, this adaptation cannot be visualised.
Table 7: Evaluation of cases with 2000 MCS iterations ($O_{PD}$ in days, $O_{C}$ in $10^{6}$, $O_{N}$ in nuisance level [0,10]). Where the MOO case considers $w^{\prime}_{1,PD}=0.70,w^{\prime}_{2,C}=0.15,w^{\prime}_{1,N}=0.15$.
Case P50 P80 P90
$O_{PD}$ $O_{C}$ $O_{N}$ $O_{PD}$ $O_{C}$ $O_{N}$ $O_{PD}$ $O_{C}$ $O_{N}$
SOO $O_{C}$ (C) 1463 0.26 1.42 1468 0.40 1.98 1471 0.48 2.45
SOO $O_{N}$ (N) 1459 1.38 0.36 1464 1.79 0.72 1466 1.86 0.98
MOO (M) 1473 0.36 0.62 1479 0.53 1.17 1482 0.62 1.51
abs. diff. (C - N) 4 -1.12 1.06 4 -1.39 1.26 5 -1.38 1.47
abs. diff. (C - M) -10 -0.10 0.80 -11 -0.13 0.81 -11 -0.14 0.94
abs. diff. (N - M) -14 1.02 -0.26 -15 1.26 -0.45 -16 1.24 -0.53
rel. diff. (C - N) 0.3 -430.8 74.6 0.3 -347.5 63.6 0.3 -287.5 60.0
rel. diff. (C - M) -0.7 -38.5 56.3 -0.7 -32.5 40.9 -0.7 -29.2 38.4
rel. diff. (N - M) -1.0 73.9 -72.2 -1.0 70.4 -62.5 -1.1 66.7 -54.1
The following conclusions can be made from the comparative evaluation outcome:
  1. 1.
    Comparing the three cases it becomes evident that for all percentiles the relative difference in the project duration ($O_{PD}$) is very little (1% difference between cases) suggesting that all cases perform well with optimising towards meeting the target duration. However, the SOO $O_{C}$ and SOO $O_{N}$ are only able to achieve this by the implemented contractual penalty (considered as a trade-off technique). The outcome is thus highly influenced by the defined penalty, limiting the reflection of actual project reality but instead only the contractual ”pseudo-reality’. The MOO approach achieves similar results given the desired preference of all involved stakeholders allowing for an independent representation towards the project duration and control cost, overcoming the need for monetisation techniques. Moreover, these monetisation techniques are limited with regard to ”soft” objectives (as traffic nuisance) where direct cost relation and quantification is not possible.
  2. 2.
    As expected, larger relative differences in case of project planning costs and traffic nuisance are visible. The SOO $O_{C}$ results in an average of 33% lower cost compared to the MOO approach. The SOO $O_{N}$ results in an average of 63% lower traffic nuisance compared to the MOO approach. However, when considering a stakeholder-oriented behaviour, it becomes evident that the two SOO cases reflects a narrow and one-sided view not accounting for other set of project goals. With the SOO approaches no synthesis that benefits the overall project outcome is achieved, revealing a significant drawback. The MOO outcome however resulting in acceptable levels of all respective objectives and the overall highest aggregated preference for the respective percentiles. This clearly demonstrates that the synthesis towards all project objectives can result in overall group satisfaction. Similar to the strategic planning application, utilising the MOO approach offers a more balanced perspective by taking into account multiple objectives, including the inherent uncertainties and stochastic nature of the project leading to overall favourable project realisations. Similar to the prior example it becomes evident in Figure 4 that with increasing percentile, the preference decreases concerning their respective objectives. As expected, this trend indicates that decision-makers face lower preference outcomes when targeting higher objective probabilities.
  3. 3.
    The criticality index (see Figure 5 and Figure 6) aids decision-making by clarifying the utilisation of specific control measures and their combinations, thereby identifying the effectiveness of certain control measures and combinations given the underlying uncertainties and stochastic factors. Comparing the three cases, the criticality index differs significantly where certain measures seem to have clear positive impact towards certain objectives. For example measure 18 is utilised in 80% of the iterations in the SOO $O_{C}$ however shows little impact for the SOO $O_{N}$ or MOO case. The results suggest that some corrective measures (i.e. with low occurrence frequency, e.g.: 14, 15, 16, 17) can be ignored, while others should be prioritised. The frequency of measures given SOO $O_{N}$ and SOO $O_{N}$ do not consider the impact a certain measure has towards other objectives, neglecting associative decision-making. The MOO approach reflects the frequency of measures that have the highest aggregated group preference towards all objectives revealing their actual impact. However, to arrive at effective project control strategies, it is important not to rely solely on the most critical corrective measure as a standalone solution. Instead, these measures should be combined to form a comprehensive mitigation strategy selected by the project manager. To this extent, the criticality index of measure combinations gives insight into the correlated impact of measures. The likelihood of timely project completion should then be verified through a simulation allocating these specific measures. Similar critically analysis can be performed for the project activities and project paths (see Kammouh et al. (2021)).
These observations provide project managers with actionable insights, enabling them to define the most effective project control measures given all the project objectives. A dynamic adaption and verification thought the project stages is enabled with the ability to adapt the set of feasible measures.
Figure 5: Criticality index of mitigation measures for all three optimisation scenarios.
Figure 6: Criticality index of mitigation measures combinations for all three optimisation scenarios.
5 Discussion & next steps for further development
The two example application demonstrate the advances and added benefit of the introduced Odycon methodology. Facilitating associative decision-making in uncertain projects come with limitations towards the technical and social integration. The challenge lies in the combined complexity of synthesising the concurrent interests of all stakeholders while managing intricate technical interdependencies effectively. The methodology is built for computer-aided and data-driven decision-making, relying heavily on the availability and accessibility of large project performance information and stakeholder preferences mapping. For example, estimates for activity durations, technical performance data of engineering assets, occurrence probabilities of risks, and data on external environmental influences significantly impact the interdependency and workability of a project. To achieve reproducible and consistent outputs, the variability in data quality from different sources must be minimal. Moreover, accurately representing stakeholder preferences and their relative individual weights significantly impacts Odycon’s outcomes. Consequently, well-informed project decision-making is fundamentally limited by the availability and quality of both preference and performance information.
While the introduced Odycon methodology is adaptable to various project specifics and constraints, the inflexibility of existing contracts may hinder the accommodation of dynamic changes. Legal constraints and compliance issues could limit the effectiveness of this open and cooperative decision-support methodology, impeding adaptive and associative project management.
Towards that extent, the following further research and development is recommended:
  1. 1.
    As demonstrated in the example applications, the results depends on a good reflection of human objectives and their preferences & weights (social integration). Further research is needed towards establishing a Structured Stakeholder Judgement model to improve and refine the preference elicitation to estimating the preference functions and weight distributions. This type of model could improve cooperative decision-making, promoting a collective mindset shift among all involved parties to fully utilize the benefits of the Odycon methods for the advantage of all project associated stakeholders. As a first step, the so-called choice-based conjoint analysis method can be introduced to generate an initial estimate for this.
  2. 2.
    To improve the accurate reflection of uncertainty (technical integration) the implementation of Structured Expert Judgment (e.g. Cooke’s model) should be considered when decision-makers are confronted with insufficient data availability (Leontaris et al., 2019). In case of large sets of available field-data, the use of real-time data methods like non-parametric Bayesian Networks could be integrated to model uncertainty without relying on predefined distributions (Leontaris et al., 2018).
  3. 3.
    The example applications considered Beta-PERT modelling, neglecting other simulation-based models. To improve the simulation of the network, the impact of models that consider the integration of different types of activity links and automated changes in the network structure should be explored (Wang, 2005).
  4. 4.
    Constraints regarding resources limitation (e.g., personnel, material, assets) are currently neglected. However, for more realistic applications, the availability of certain resource pools during the project execution should be incorporated to give project managers a more accurate representation of feasible project planning and control strategies.
  5. 5.
    Both postmortem analysis and real life projects research can further support the effect of using Odycon and add confidence to contractor-client relationships so that the current rigid forms of contract may become more open.
  6. 6.
    Further research can be done to use Odycon in stalemate projects as a transparent and objective mediator to ’confront project conflicts into yes’.
  7. 7.
    The Odycon methodology can be further extended for project portfolio optimisation to reduce inefficiencies.
6 Conclusion
The increasing complexity of projects comes with the need for stakeholder and goal-oriented decision-making, and is addressed by following a systems-oriented approach.
This paper presents Odycon, a pure a-priori stochastic simulation & optimisation methodology integrating the capability of the project (technical domain), the human goal-oriented behaviour (human domain), and the association of stakeholder-oriented behaviour (social domain). To this extent, the IMAP optimisation method is integrated into a probabilistic MCS, accurately reflecting project uncertainties while enabling best-fit for common-purpose project management. The advances towards strategic planning and dynamic control were demonstrated with two distinct examples demonstrating the utilisation during up-front planning and on-the-run project control. The following general conclusions can be drawn.
First, the complex reality of project management is addressed by identification of the most effective and desired set of project planning & control variables given the uncertainties and stochastic factors while considering overall group satisfaction. The interplay with stakeholder-oriented concurrent objectives is addressed by the integration of the ’associative preference domain’, enabling an optimal solution that best fits the common purpose, using a-priori optimization. This approach supports satisfactory project planning and control, allowing stakeholders to meet their individual needs through collective performance. The model results of both demonstrative applications show that joint project success, in alignment with the defined common purpose, will increase when individual stakeholders relinquish their pure self-interest, ultimately enabling an optimal solution that benefits the whole group. Here, Odycon aligns with previous empirical results of management studies, as mentioned in the introduction. The developed methodology allows for the integration and combination of variables from different project phases, as well as flexible adaptation and development tailored to project and stakeholder specific needs.
Second, Odycon opens the possibility to reflect all types of project goals, with no need for cost or resource trade-off techniques, thus providing flexibility and better insight into different objectives and system behaviour. It enables adaption towards project specific circumstances where numerous important factors may not be covered by contractual specifications or budget, leading to a decline in quality under the conventional management practice. This consideration is essential for enhancing project planning and control strategies while maintaining project integrity in terms of quality, budget, and timeline. While the budget of a project is determined by its financial limits, other objectives describing the quality or impact of a project should be evaluated based on their degree of ’satisfaction,’ which reflects the utility or value they offer.
Third, Odycon provides clear guidance in budget and resource allocation before and during project execution, making it an essential tool to support decision-making. Project managers and associated stakeholders are able to jointly create actionable insights into the system behaviour and decision-making impact of planning and/or control variables given the underlying uncertainties a project faces. This enables to jointly allocate planning and control budgets more effectively towards solutions that are both feasible and desired.
Fourth, Odycon enables well-informed and deliberative decision-making by removing bias from the decision-making process. Manual trial-and-error approaches are time-consuming and inefficient, relying on ’gut instinct’ and human judgment on uncertainties, which can be biased and limited. In contrast, Odycon’s computer-aided decision-support framework uses simulation and optimisation models to explore numerous solutions systematically, reducing human biases, enhancing efficiency, and acknowledging solution spaces no longer conceivable for humans. Odycon thus facilitates a shift within complex project management from unsubstantiated trial-and-error to transparent and data-driven decision modelling, optimizing over large numbers of variables and objectives, within a multidimensional solution space.
With this methodology, Odycon takes a next step in computer-aided data-driven decision-making for project and operations management through significantly improving the efficiency and effectiveness of decision-making, thereby enabling an associative ideal within reach.
Disclosure statement
The authors report that there are no competing interests to declare.
Data availability statement
The two demonstrative applications for strategic project planning and dynamic project control are available via a GitHub repository online in accordance with confidentiality agreements. The repository is located here: https://github.com/Boskalis-python/ODYCON. Note, for preference aggregation as part of the optimisation, the following algorithm is used: https://github.com/Boskalis-python/a-fine-aggregator.
Appendix A Preference score aggregation
The following describes the aggregation algorithm for use in preference-based design and decision-making. In short, the following two starting principles apply to this algorithm (see Barzilai (2022) or Wolfert (2023)):
  1. 1.
    it should reflect relative scoring as encountered in actual design and decision-making practice.
  2. 2.
    it should adhere to the governing mathematics in a one-dimensional affine space, which is the mathematical model applicable to preference score(s).
The algorithm therefore consists of two operations: (1) normalising the preference scores of all alternatives per criterion, and (2) finding the representative aggregated preference score $P^{*}$ for each alternative using the weighted least squares method. These two operations are further elaborated mathematically below.
A.1 Normalisation
For normalisation, the standard score (z-score) method is used. This yields a normalisation that preserves information about the population of preference scores and reads as follows:
$z_{i,j}=\frac{p_{i,j}-\mu_{j}}{\sigma_{j}}\textrm{ for }i=1,2,...,I\textrm{; }j=1,2,...,J$ (15)
Here $z_{i,j}$ is the normalised score of alternative $i$ for criterion $j$; $p_{i,j}$ is the preference score of alternative $i$ for criterion $j$; $\mu_{j}$ is the mean of all preference scores $p$ for criterion $j$; $\sigma_{j}$ is the standard deviation of all preference scores $p$ for criterion $j$. By performing this normalisation for all criteria $J$, the preference scores are transformed to a single scale with the same properties ($\mu_{J}=0$, $\sigma_{J}=1$).
A.2 Weighted least squares
Since all $z_{i,j}$ scores are now on a single scale, it is possible to compare all normalised scores per alternative with each other. To find the representative aggregated preference score of an alternative that provides a best fit of all (weighted and relative) scores for each criterion, a minimisation of the weighted least squares difference between this aggregated score and each of the (normalised) individual scores on all criteria is applied. This is expressed mathematically as follows:
Minimise $\textit{Minimise}\text{\ }S_{i}=\sum_{j=1}^{J}w_{j}*(z_{i,j}-P_{i}^{*})^{2}$ (16)
Note that since the search is for a single representative aggregated preference score, the model function $f(x_{i,j},\beta_{i})$ from the classical weighted least square method is replaced by $P_{i}^{*}$. The solution to this minimisation can be found by differentiating with respect to $P_{i}^{*}$ and equating it to zero. Since $\sum_{j=1}^{J}w_{j}=1$, this results in the following analytical expression for the representative aggregated preference score:
$P_{i}^{*}=\sum_{j=1}^{J}w_{j}*z_{i,j}$ (17)
Appendix B Dynamic control application data
Figure 7: Simplified project planning and SAAone’s construction sequence (logical network model) represented as a Gantt chart, a (deterministic) critical path is represented in red. This was also assumed as the base project planning in Kammouh et al. (2022).
Table 8: Dynamic control application: Project Activities.
ID Activity description Activity duration (days) Predecessors
$a_{i}$ $m_{i}$ $b_{i}$
1 Contract date 0 0 0 0
2 Financial Close 0 0 0 0
3 Design 819 920 1435 1
4 Acquiring the certificate of commencement 105 130 194 1
5 Commencement certificate is acquired 0 0 0 2,4
6 Date of commencement 0 0 0 5
7 Maintain existing road assets A9 / A1 / A6 976 1284 1836 6
8 Conditioning, Cables and Conducts, permits 168 200 268 6
9 Preload 324 395 525 6
10 Constructing a new Aqueduct 200 260 341 6
11 Constructing a Canal bridge 285 335 492 6
12 Construction works in the southern A1 lane 113 128 189 9,10,11
13 Commissioning of the southern A1 new lane 0 0 0 12
14 Producing parts of Railbridge part 1 223 251 366 6
15 Producing parts of Railbridge part 2 194 220 350 14
16 Assembling a railway bridge on location 559 674 971 14
17 Moving Railway Bridge in place during Train Free Period 0 0 0 16
18 Road works northern A1 lane 109 130 191 17
19 Commissioning of the Northern A1 new lane 0 0 0 18
20 Road and construction works new part junction Diemen 477 530 848 13
21 Build new viaducts A6 304 400 532 6
22 Build second Hollandse bridge 286 340 459 6
23 Road and construction works junction Muiderberg 716 930 1237 6
24 Road works eastern part A6 90 100 130 21,22
25 A6 East ready 0 0 0 24
26 Reconstruction western part A6 324 400 532 25
27 Commissioning A6 0 0 0 18,23,26
28 Road works existing part Diemen junction 71 90 130 19,20
29 Request Availability Certificate 0 0 0 28
30 Assess and obtain Availability Certificate 16 20 27 29
31 Demolition old A1 (part 1) 54 61 91 30
32 Demolition old A1 (part 2) 23 30 48 31
33 Greenery for old A1 77 90 129 31
34 Applications and obtaining partial completion certificates 104 120 161 30
35 Request Completion Certificate 0 0 0 33
36 Obtaining the Certificate of Completion 17 20 27 35
37 Scheduled Completion Date 0 0 0 36
Table 9: Strategic control application: Shared uncertainty factors.
Shared uncertainty (days)
ID Shared uncertainty factor $a_{i}$ $m_{i}$ $b_{i}$ Relations with activities
1 Weather condition 1 -45 0 72 10,11
2 Soil composition -50 0 100 $21,22,23,7$
3 Crew performance -10 0 50 12,23
4 Soil composition -45 0 110 20,26
5 Equipment availability 1 -20 0 100 15,16
6 Site availability -5 0 100 16,20
7 Procurement, fabrication or assembly -1 0 55 7,20
8 Project control and management -20 0 50 8,9
9 Design or documentation accuracy -5 0 15 32,33
10 Owner-driven changes 0 0 45 18,20
11 Issues with contractor -20 0 50 3,4
12 Issues with supplier -20 0 100 7,14
13 Equipment availability 2 -80 0 90 7,16
14 Weather condition 2 -140 0 100 7,23
Table 10: Dynamic control application: Risk events.
ID Risk event description Risk duration (days) Relation $p_{e}\quad$
$a_{i}$ $m_{i}$ $b_{i}$
1 Preliminary design rejection, including extra scope of works 96 105 119 3 0.20
2 EDP audit failure 13 14 15 4 0.05
3 Condition deviates from plan 63 70 78 7 0.15
4 Unexpected gas conducts 35 35 41 8 0.20
5 Lower consolidation rate than calculated for 34 35 40 9 0.10
6 Piling machines break down 14 14 15 10 0.10
7 Late delivery of prefab elements 19 21 25 11 0.20
8 Dynamic traffic management equipment/software not functioning 20 21 22 12 0.25
9 Production equipment failure 20 21 23 14 0.05
10 Construction site subsides 13 14 15 15 0.05
11 Ancillary equipment failure 33 35 41 16 0.10
12 Dynamic traffic management equipment/software not functioning 20 21 21 18 0.25
13 Discovery of polluted soil 13 14 14 20 0.05
14 Concrete casting failure 13 14 14 21 0.05
15 Main pillar subsides 65 70 71 22 0.02
16 Discovery of polluted soil 25 28 32 23 0.05
17 Insufficient quality of base layer 39 42 47 26 0.02
18 Discovery of asphalt with too high PAK percentage 13 14 17 28 0.05
19 Additional scope of work (miscellaneous) 130 140 160 30 0.01
Table 11: Dynamic control application: Corrective measures (variables).
Capacity [days] Cost [€] $(\eta=0.5)$ Nuisance
ID Mitigation description $a_{i}$ $m_{i}$ $b_{i}$ $a_{i}$ $m_{i}$ $b_{i}$ $a_{i}$ $m_{i}$ $b_{i}$ Relation
$x_{1}$ Extra engineering design office personnel 99 101 101 118k 120k 120k 0.00 0.00 0.00 3
$x_{2}$ Extra software design capacity 14 14 14 30k 30k 30k 0.00 0.00 0.00 4
$x_{3}$ Extra maintenance engineers 103 127 127 136k 150k 150k 0.91 1.00 1.00 7
$x_{4}$ Extra administrators for permitting 43 51 57 44k 48k 50k 0.00 0.00 0.00 8
$x_{5}$ Applying extra preloading material 41 51 51 677k 750k 750k 5.42 6.00 6.00 9
$x_{6}$ Adding extra onsite construction flow 92 101 107 190k 200k 205k 9.52 10.00 10.00 10
$x_{7}$ Extra prefab construction capacity 117 127 129 144k 150k 151k 0.96 1.00 1.01 11
$x_{8}$ Extra M&E engineers 51 51 51 60k 60k 60k 3.00 3.00 3.00 12
$x_{9}$ Extra welding equipment and personnel 53 64 64 90k 100k 100k 5.45 6.00 6.00 14
$x_{10}$ Extra temporary soil measures 45 51 53 235k 250k 254k 5.65 6.00 6.12 15
$x_{11}$ Ancillary on standby 201 203 222 199k 200k 209k 4.98 5.00 5.24 16
$x_{12}$ Extra M&E engineers 14 14 14 30k 30k 30k 3.00 3.00 3.00 18
$x_{13}$ Extra excavation capacity 96 101 101 121k 125k 125k 9.71 10.00 10.00 20
$x_{14}$ Extra concrete workers/carpenters 82 101 107 67k 75k 77k 5.42 6.00 6.17 21
$x_{15}$ Temporary ancillary construction and rework 70 76 84 1,442k 1,500k 1,576k 7.69 8.00 8.41 22
$x_{16}$ Extra excavation capacity 60 76 82 134k 150k 155k 7.18 8.00 8.31 23
$x_{17}$ Extra asphalt equipment and personnel 101 101 107 200k 200k 205k 8.00 8.00 8.23 26
$x_{18}$ Extra removal works 43 51 53 69k 75k 76k 9.23 10.00 10.00 28
$x_{19}$ Extra equipment and personnel 10 14 16 214k 250k 267k 6.86 8.00 8.57 30
$x_{20}$ Applying extra preloading material during the night 41 51 51 1,084k 1,200k 1,200k 0.90 1.00 1.00 9
$x_{21}$ Automation of M&E workflows 30 30 30 120k 120k 120k 0.00 0.00 0.00 12
$x_{22}$ Welding operation during the night 45 53 53 166,k 180k 180k 1.85 2.00 2.00 14
$x_{23}$ Extra temporary soil measures during the night 32 40 42 315k 350k 358k 0.90 1.00 1.03 15
$x_{24}$ Night shifts for concrete work 67 80 91 91k 100k 106k 3.68 4.00 4.28 21
$x_{25}$ Excavation during the night 40 56 62 257k 300k 316k 0.86 1.00 1.05 23
$x_{26}$ Removal works during the night 33 41 43 81k 90k 92k 0.90 1.00 1.02 28
$x_{27}$ Lane by lane asphalt work 30 30 30 430k 430k 430k 3.00 3.00 3.00 26

### v2

@@
 Abstract
-Strategic project planning and dynamic control is essential for ensuring complex projects to be executed best-fit for common purpose.
-When best-for-project strategies are no longer conceivable for humans, there is a need for effective and efficient computer-aided decision support.
-To this end, standard simulation-driven evaluation and a-posteriori decision-making is often used instead of a combined simulation and optimisation approach that a-priori integrates the stakeholders goal orientation and risk management into a best-fit design solution.
-However, recently developed state-of-the-art planning and control methodologies that do already incorporate this integration still lack a stochastic representation and/or an associative multi-objective approach.
-This paper presents a new project management methodology, called Odycon (Open design and dynamic control), which dissolves the aforementioned shortcomings.
-To enable this, a generic mathematical statement is first formulated for project planning and dynamic control uniting technical (logistical) capabilities, human goal-oriented behaviour, and stakeholders interests.
-Then, both Monte-Carlo simulation (MCS) and the Integrative Maximisation of Aggregated Preferences (IMAP) optimisation method are combined, resulting in a best-fit strategic planning and dynamic control methodology.
-Odycon’s use and added value are demonstrated for two applications: (1) a pure strategic planning of an offshore wind installation project, and (2) a pure dynamic control of a highway infrastructure construction project.
-Both applications demonstrate Odycon’s advances in concurrent and associative design and decision-making, offering best-fit for common purpose synthesis for different complex project phases.
+Strategic project planning and dynamic control are essential to ensure that complex projects are executed in a best-fit-for-common-purpose manner.
+When such strategies become inconceivable for humans, effective and efficient computer-aided decision support becomes crucial.
+Standard simulation-driven evaluation and a-posteriori decision-making are often used, rather than a combined simulation and optimisation approach that a-priori integrates the stakeholder interests and risk management.
+However, current planning and control methodologies often lack stochastic representation and/or associative multi-objective optimisation approaches within socio-technical reach.
+This paper introduces Odycon (Open design and dynamic control), a new purpose-driven project management by design methodology that resolves these shortcomings.
+A generic mathematical project planning and control statement is developed to integrate and represent both stakeholders preferences (human domain) and project system performances (technical domain) for effective project planning and dynamic control.
+Odycon integrates Monte-Carlo simulation (MCS) with the Integrative Maximisation of Aggregated Preferences (IMAP) optimisation method to develop a best-fit strategic plan and dynamic control measures.
+Its value is demonstrated through applications in offshore wind installation and a highway infrastructure projects, showcasing advances in associative design and decision-making, and aiming for a best-fit-for-common purpose synthesis across different complex project phases.
 keywords:
-Project management , Project scheduling , Project control , Risk management , Multi-objective optimisation , Monte Carlo simulation , Preference function modelling
-journal: arXiv
+Project management , Project planning & scheduling , Stochastic Project control , Risk management , Multi-objective optimisation , Preference & Performance-based optimization , Monte Carlo simulation , Preference function modelling
 […]
@@
 [b]organization=Corporate R&D, Boskalis,addressline=Rosmolenweg 20, city=Papendrecht, postcode=3356 LK, state=Zuid-Holland, country=The Netherlands
+[…]
+[c]organization=Department Algorithmics, Faculty of Mathematics and Computer Science, Delft University of Technology,addressline=Van Mourik Broekmanweg 6, city=Delft, postcode=2628 XE, state=Zuid-Holland, country=The Netherlands
 1 Introduction
-Successful project delivery is traditionally attempted through project planning and control, answering what is going to be done when and how much does it cost.
-Planning is referred to the process of deciding how to execute a project, while project control reflects the process of ensuring execution as planned.
-Establishing a robust and reliable project schedule is fundamental for preliminary cost and time estimates, as well as for effective and efficient project management (Smith et al., 2014).
-However, with increasing size and degree of uncertainty, initial estimates of costs and timelines tend to be fraught with inherent risks and uncertainty that cannot be completely eliminated.
-These uncertainties often results in discrepancies between early predictions and stakeholder objectives, complicating the decision-making process and may result in unmet expectations.
-Initial project planning and execution control is often approached by individual optimisation or simulation methods, resulting in a one-sided point of view and neglecting the multifaceted nature of projects where the technical capabilities, the human goal-oriented behaviour, and the stakeholders desirability are interconnected.
-It requires a systems-oriented approach to acknowledge these complexities and recognise the need for stakeholder-oriented decision-making to answer what should be accomplished why and how (Reschke and Schelle, 2013).
-The role of well-informed human decision-making is crucial in project management, where decision-makers must understand the rationale behind their chosen project planning or control strategies towards risk management, maximum effectiveness, and ultimately stakeholder satisfaction.
-When evaluating the overall project’s success, it becomes clear that factors beyond just cost and timely completion play a significant role (Van Gunsteren, 2011).
-Trade-offs between quality, costs, and schedule are a constant reality in construction projects and are typically determined by the involved stakeholders.
-These trade-offs are reflected by the socio-technical interplay between the stakeholders preference (human domain) and the capability of the project (technical domain), see Van Heukelum et al.
+In today’s complex engineering project management landscape, wicked problems often cause projects to derail.
+In fact, projects don’t go wrong — they start wrong (Flyvbjerg, 2024).
+And even when they start right, they can still lose direction.
+Advanced design and project management practices may still hit impasses because project managers often rely on retrospective analyses (such as those by Flyvbjerg (2024) or others like Yang et al.
+(2022)) and traditional decision-support methods, which offer little guidance for a constructive way forward.
+These challenges become particularly acute when decision-makers fail to recognize that their problems are part of a larger, interconnected whole — and that mutual interconnection and collaboration are key to unlocking complex projects that achieve a best-fit for a common purpose.
+This calls for moving beyond the traditional notion of fitness for purpose (total quality), which emphasizes meeting predefined requirements and delivering technically effective results.
+Instead, we must adopt a broader, value-driven perspective — one that focuses on what is socially desirable and physically feasible from a shared purpose.
+Best-fit for common purpose integrates stakeholder interests, dynamic context, and collective value creation within socio-physical reach.
+Moreover, what is needed is a proactive threefold approach to confront project complexity — ensuring that projects start, proceed, and finish successfully: (1) Agreeing First on a best-fit scope, (2) Acting Feasibly with an optimal project plan, and (3) Adapting Flexibly through the best set of mitigation measures.
+Establishing a robust and socio-technically feasible project plan is fundamental to effective and efficient project management (Smith et al., 2014).
+However, as project size and uncertainty increase, initial estimates regarding scope, cost, and timelines are often exposed to inherent risks and uncertainties that cannot be entirely eliminated.
+Project success is therefore achieved through dynamic adaptation to emerging challenges—often not specified in a static contract—and through collaborative engagement among stakeholders to develop associative solutions.
+Focusing solely on the final scope as defined in the contract, insisting on creating a fixed plan and adhering rigidly to it, is ultimately naïve and pseudo-realistic in the face of real-world project dynamics.
+This ultimately calls for an actionable, open design and systems-oriented approach, supported by transparent decision-support tools that foster both deliberate (slow) and socially informed thinking.
+As such, successful project delivery hinges on purpose-driven and optimal strategic planning and dynamic control, transforming complexity into outcomes that best fit within socio-physical reach.
+Planning refers to the process of reaching agreement first in order to act feasibly, while project control ensures execution according to plan and the ability to adapt flexibly when needed on the run.
+Despite this reality, traditional project management in construction often remains heavily focused on cost and schedule, with stakeholders primarily considering these from individual perspectives.
+This narrow focus neglects the interplay of stakeholder-oriented, concurrent objectives.
+Joint project success can only be enhanced when stakeholders are willing to look beyond self-interest and engage in a process of mutual alignment.
+This requires stakeholders to let go of their individually defined objectives in favour of a synthesis solution that best fits the common purpose — one that ultimately outperforms single-sided or even compromise-based project outcomes (Wolfert, 2023; Scharmer, 2016; Glasl, 1998; Van den Doel and Van Velthoven, 1993).
+By integrating diverse stakeholder interests and accounting for project capabilities in the decision-making process, a more holistic outcome—referred to as the best-fit for common purpose—can be achieved, rather than settling for suboptimal compromises (Zhilyaev et al., 2022).
+This best-fit design emerges from the socio-technical interplay between stakeholder preferences (the human domain—what they want) and system performance capabilities (the technical domain—what it can deliver), as developed and demonstrated by Van Heukelum et al.
 (2024).
-However, the traditional best practice recommends solely adhering to the contract and making a plan without considering the actual project reality, often following the rule to make a plan and stick to it.
-This practice can be conceived as ’pseudo-reality’, as project success is created by the dynamic adaptation to issues that arise (most often not defined in a contract) and dynamic interaction between stakeholders towards associative solutions.
-Despite this reality, the traditional managerial focus in construction projects often leans heavily towards costs and schedule and their individual interest, thus neglecting the interplay with stakeholder-oriented concurrent objectives in the process.
-When stakeholders are willing to go beyond their pure self-interest, joint project success can be enhanced.
-Stakeholders do have to relinquish their identified individual goals, ultimately enabling a synthesis solution that best fits the common purpose and outperforms single-sided project or even compromise solutions (Glasl, 1998; Scharmer, 2016; Van den Doel and Van Velthoven, 1993; Wolfert, 2023).
-Traditional planning & control methods lack to explicitly express the relationship and influence of all participating stakeholder interests towards the decision-making process in project or operations management (Blanchard and Fabrycky, 2011; Del Pico, 2023; Slack et al., 2010).
-Usually (both in science and in industry), these methods are simulation and analysis oriented (a-posteriori) and do not take an (a-priori) design approach to effectively and efficiently generate a best-fit management solution.
-However, state-of-the-art methods that are optimisation-based often do not reflect a best-fit for common purpose consideration.
-These are most often being limited to a single objective optimisation (SOO) approach, relying on monetisation to reflect multiple objectives and neglecting the complexity of stakeholder satisfactory project management.
-While the benefit of multi-objective optimisation (MOO) in project management has been acknowledged (Guo and Zhang, 2022), there has been no methods developed towards adapting MOO in strategic project planning and dynamic project control integrating probabilistic risk management (Kammouh et al., 2022).
-Well-informed decision-making is facilitated by removing overconfidence bias through the use of structured and data-driven decision-making processes (Kahneman, 2011).
-Considering the inclusion of the goal-orientation of all involved stakeholders and their desired expectations, the increasing complexity of construction projects necessitates effective and efficient computer-aided decision support for stakeholder-oriented project management.
-To address the lack of consideration for human control and adaptive behaviour in standard probabilistic scheduling methods using Monte-Carlo simulation (MCS) (Del Pico, 2023), research by Kammouh et al.
-(2022, 2021) has proposed the mitigation controller (MitC), a current state-of-the art project control software.
-It automates the project manager’s selective nature to only apply the most effective and efficient control measures on-the-run, based on the unique circumstances of each possible project realisation.
-The MitC, is designed as a SOO control tool, optimising for cost-effectiveness given a project delivery date (target constraint).
-As a result, there is no understanding of how the most optimal set of control measures impacts other project objectives from relevant decision-makers.
-While other objectives can be assessed independently, there is no guarantee that an optimal project planning and control strategy is derived that represents a best-fit for common purpose synthesis.
-Moreover, with increasing project complexity, the impact of individual planning & control variables related to all stakeholder objectives becomes cumbersome to be evaluated in a manual trial-and-error approach.
-Associative project planning and control enabling a stakeholder-oriented behaviour can be introduced by utilising the advances developed in the Open Design Systems (Odesys) methodology (Wolfert, 2023), that enables the association of the individual stakeholder interests (desirability) into the technical design optimisation (capability) using integrative maximisation of aggregated preferences (IMAP).
-This Odesys/IMAP approach proves to be an effective methodology to arrive at best-fit for common purpose design solutions where all stakeholder interest are accounted for and translated to a common preference domain to find an overall group optimum.
-The introduced IMAP method enables iterative group design and decision-making in a deterministic manner.
-However, to effectively utilise these advancements for project planning and control, it is crucial to incorporate traditional probabilistic project planning and control methods that can precisely capture and represent the project’s schedule capabilities, such as the Program Evaluation and Review Technique (PERT).
+However, achieving such a best-fit requires more than technical alignment; it also demands a fundamental shift in how planning and execution control are approached.
+Both academic research and industry practice tend to rely on retrospective, analysis-oriented methods (a-posteriori), which offer limited support for forward-looking strategic development.
+Moreover, these methods are often driven by individual optimization or simulation-based approaches, producing one-sided solutions that overlook the multifaceted nature of real-world projects.
+In practice, technical capabilities, human goal-oriented behavior, and stakeholder desirability are deeply interconnected.
+Yet traditional planning and control approaches fail to explicitly capture the dynamic interplay of all stakeholder interests within this complex decision-making process (Blanchard and Fabrycky, 2011; Del Pico, 2023; Slack et al., 2010).
+Most existing design optimisation methods rely on single-objective optimisation (SOO), often reducing multiple project objectives to a single cost metric.
+This approach overlooks the qualitative nature of stakeholder preferences, which are essential for goal-oriented decision-making.
+While the benefits of multi-objective optimisation (MOO) in project management are increasingly recognized (Guo and Zhang, 2022), current methodologies rarely integrate MOO within strategic project planning or dynamic project control frameworks that also incorporate probabilistic risk management (Kammouh et al., 2022).
+To address the rising complexity and uncertainty in construction projects, decision support must evolve to enable computer-aided, stakeholder-oriented planning and control based on MOO.
+As Kahneman (2011) emphasizes, deliberate and data-informed processes are essential for mitigating cognitive biases such as overconfidence.
+Consequently, well-informed project decision-making requires transparent, evidence-based tools that support “slow and social thinking”—tools that not only process data but also help align diverse stakeholder objectives within feasible and resilient project solutions.
+In this paper, all the aforementioned considerations are taken into account in the development of a new open-design-driven dynamic project planning and control methodology, named Odycon (acronym: Open Design and Dynamic Control).
+Odycon builds on the core principles and optimization method of the Open Design Systems (Odesys) methodology (Wolfert, 2023).
+The Odesys integrative maximisation of aggregated preferences (IMAP) optimization method has proven effective in achieving best-fit-for-common-purpose design solutions, where all stakeholder interests are accounted for and translated into a common preference domain to identify an overall group optimum.
+Notably, Odesys/IMAP is a novel preference-based optimization methodology where, instead of directly minimizing or maximizing individual objectives, the rigorous mathematical optimization focuses solely on maximizing the aggregated preference of stakeholder objectives within physical reach.
+So far, the Odesys/IMAP method enables iterative group design and decision-making in a deterministic manner.
+However, for a project planning, it must be extended to incorporate traditional probabilistic project planning techniques that capture schedule performance, such as PERT or discrete-event simulation (DES).
+This constitutes a truly novel project management-by-design methodology.
 Additionally, these tools should employ stochastic simulation methods to reflect the inherent uncertainties and unpredictability encountered both in the planning and execution phases.
-Utilising and combining these advances in simulation and optimisation can notably reduce inefficient and misinformed decision-making in project planning and control, particularly with increasing size of projects, where it becomes virtually impossible to account for all possible scenarios.
-In this context, the focus of this paper is to introduce a methodology that integrates (1) the physical and logistical capability of the project (’network performance domain’), with (2) the human goal-oriented behaviour (’decision-maker objective domain’) and extends it towards (3) the desirability of the project stakeholders (’associative preference domain’) enabling best-fit for common purpose project management.
-This is required as the current state-of-the-art methodologies for project planning or control exhibit the above-mentioned shortcomings that make them unsuitable for delivering the necessary integrative & associative behaviour to successfully execute the project so that it best meets all stakeholder interests.
-To achieve such a best-for-project fit, the project management practice PII introduced by Van Gunsteren (2011) and based on Schön and Argyris (1996) can be further improved to a PIII level.
-This advancement goes beyond the integrative double-loop learning approach, which involves making decisions for the stakeholder or project client.
-Instead, PIII emphasizes open-loop learning by working closely in association with all project stakeholders to achieve the best possible overall project result.
-A step from stakeholder oriented, but static ’self’-planning and control, to pure dynamic ’group’-planning and organisation where an integrative maximization of associated preferences is achieved Wolfert (2023).
-Based upon the identified development gap and utilisation needs, this paper presents a new integrative stochastic simulation and multi-objective planning & control optimisation methodology, named Odycon (an acronym of open design and dynamic control).
-Odycon enables best-fit for common purpose generation of both strategic planning and/or dynamic control applications.
-Odycon is a decision-support methodology that automates the selection of planning and control variables, taking into consideration multiple stakeholder objectives and constraints.
-Building upon the MitC introduced by Kammouh et al.
-(2021) and the IMAP optimisation method introduced by Wolfert (2023) and Van Heukelum et al.
-(2024), this paper aims to advance open and dynamic project planning and control of complex construction projects, fostering multi-objective stakeholder-oriented decision-making.
-The following advances are presented:
-(1) A general framework for an integrated strategic planning and dynamic control problem, which combines probabilistic project schedule simulations with a multi-objective design optimisation.
-This allows for dynamic adaptation to project-specific needs and stakeholder objectives, enhancing associative decision-making in complex construction projects.
-(2) A direct representation between the project feasibility, project management objectives, and stakeholder decisions, achieving a unified integration of network performance, stakeholder objectives, and associative decision-making.
-(3) A pure preference-based aggregation for multi-objective decision-making for project management, which translates various objectives and resources into a common preference domain to achieve an optimised group synthesis.
-This resolves the need for monetisation of non-cost-related objectives and does not go for the lowest cost but for the highest value.
-This paper first introduces the general mathematical statement of the multi-objective project planning & control optimisation.
-Following this, the underlying framework and concept of the stochastic simulation & optimisation methodology are described, with a detailed elaboration on the design of the method, in which the IMAP optimisation is implemented.
-The methodology is then demonstrated using a pure strategic planning example (no control variables) for an offshore wind installation project, where a service provider and a marine contractor are the key decision-makers.
-In contrast to the example focusing on pure strategic planning, the paper also demonstrates a pure dynamic control example (no planning variables) for a highway infrastructure project.
-This example includes associative control by both a contractor and a highway agency, showcasing the application of the methodology for two distinct project management phases.
-Finally, recommendations for further development are provided, and the added value of the newly developed stochastic simulation and optimisation methodology is concluded.
+For project control applications he method must also be expanded to shift the focus from optimizing design variables to optimizing the allocation of control measures, using multiple project objectives to evaluate their effectiveness and efficiency.
+To achieve this, the state-of-the-art project manager-oriented mitigation controller (MitC), as developed and demonstrated by Kammouh et al.
+(2022, 2021), will be enhanced from a single-cost objective orientation to optimally facilitate stochastic, multi-objective project control.
+In doing so, Odycon constitutes a truly novel project management-by-design methodology.
+To summarize, the following Odycon advances are presented in this paper: 1.
+A general framework for integrated strategic planning and dynamic control, combining probabilistic project schedule simulations with preference-based multi-objective design optimisation.
+This enables dynamic adaptation to project-specific conditions and stakeholder objectives, thereby enhancing associative decision-making in complex construction environments.
+2.
+A direct linkage between project feasibility, project management objectives, and stakeholder decisions, resulting in a unified integration of network performance, stakeholder values, and associative planning and control processes.
+3.
+A pure preference-based aggregation mechanism for multi-objective decision-making in project management.
+This translates diverse objectives and constraints into a shared preference domain, enabling an optimised group synthesis without relying on monetisation.
+Rather than selecting the lowest cost solution, the approach supports identifying the solution with the highest combined value for both project outcomes and people—that is, a best-fit for common purpose plan and set of control measures.
+4.
+A methodological advancement from PII to PIII in project management practice.
+Building on the PII model for complex construction projects introduced by Van Gunsteren (2011), itself rooted in double-loop learning as described by Schön and Argyris (1996), this paper proposes an evolution toward PIII—Best Practice for Collaborative Projects.
+Unlike PII, which focuses on making improved decisions for internal stakeholders via integrative reflection and learning, PIII promotes associative open-loop learning and co-creation with all relevant stakeholders.
+This fosters dynamic group-based planning and control, where integrative maximisation of associated preferences replaces static, self-oriented decision-making.
+In doing so, it enables a genuinely collaborative approach to achieving the best possible overall outcome for both project and people (Wolfert, 2023).
+Combining the advances of MitC and IMAP enables a comprehensive representation of all stakeholder preferences and objectives in project planning and control, while simultaneously accounting for socio-physical constraints.
+This significantly reduces the risk of inefficient or misinformed decision-making—especially as project size and complexity increase, making it virtually impossible to manually consider all potential scenarios.
+Odycon supports the generation of best-fit-for-common-purpose solutions for both strategic planning and dynamic control applications.
+As a decision-support methodology, Odycon automates the selection of planning and control variables by integrating multiple stakeholder objectives and constraints into a unified stochastic simulation and optimisation process.
+This paper is structured as follows.
+It begins by introducing the general mathematical formulation of Odycon’s multi-objective project planning and control optimisation approach.
+Next, the underlying framework and core concepts of the stochastic simulation and optimisation methodology are explained, including a detailed description of the method’s design, within which the IMAP optimisation is embedded.
+The methodology is first demonstrated using a strategic planning case (with no control variables) involving an offshore wind installation project.
+In this scenario, a service provider and a marine contractor act as the primary decision-makers and are able to agree on a shared scope, resulting in an actionable and feasible optimal plan.
+Following this, a dynamic control case (with no planning variables) is presented, applied to a highway infrastructure project.
+This demonstrator features associative mitigation control by both a contractor and a highway agency, illustrating how to arrive at a best set of control measures to enable flexible adaptation across two distinct project management phases.
+The paper concludes with recommendations for further development and reflects on the added value of the newly developed stochastic simulation and optimisation methodology.
 2 Mathematical formulation of the multi-objective optimisation
@@
   • 3.
-    $P_{k,i}(O_{i}(\mathcal{N}(\mathbf{x},\mathbf{y}),F_{1,2,...,J}(\mathbf{x},\mathbf{y}))$: Preference functions that describe the preference stakeholder $k$ has towards objective function $i$, which are functions of different performance functions and dependent on planning and/or control variables and physical variables ( $i\leq k$ and $K$ is the maximum number of stakeholders).
+    $P_{k,i}(O_{i}(\mathcal{N}(\mathbf{x},\mathbf{y}),F_{1,2,...,J}(\mathbf{x},\mathbf{y}))$: Preference functions that describe the preference stakeholder $k$ has towards objective function $i$, which are functions of different performance functions and dependent on planning and/or control variables and (physical) variables ( $i\leq k$ and $K$ is the maximum number of stakeholders).
   • 4.
@@
   • 5.
+    $\mathcal{N}(\mathbf{x},\mathbf{y})$: Performance function that describes the project network, depending on planning and/or control variables $\mathbf{x}$ and (physical) variables $\mathbf{y}$.
+Here $\mathcal{N}$ is a logical representation of a project planning network, composed of nodes (activities) and edges (logical links) with their respective properties.
+  • 6.
     $\mathcal{N}(\mathbf{x},\mathbf{y})$: Performance function that describes the project network, depending on planning and/or control variables $\mathbf{x}$ and physical variables $\mathbf{y}$.
-Here $\mathcal{N}$ is a logical representation of a physical project planning network, composed of nodes (activities) and edges (logical links).
-  • 6.
+Here, $\mathcal{N}$ is a parametric graph representation of a project planning network, composed of nodes (activities) $V$ and edges (logical links) $E(\mathbf{x},\mathbf{y})$.
+Each node $v\in V$ has attributes such as duration $D_{v}(\mathbf{x},\mathbf{y})$, risk $R_{v}(\mathbf{x},\mathbf{y})$, and cost $C_{v}(\mathbf{x},\mathbf{y})$, while each edge $e\in E(\mathbf{x},\mathbf{y})$ has attributes such as weight $W_{e}(\mathbf{x},\mathbf{y})$ and capacity $C_{e}(\mathbf{x},\mathbf{y})$.
+  • 7.
     $F_{1,2,...,J}(\mathbf{x},\mathbf{y})$: Performance functions that describe the object behaviour, depending on the planning and/or control variables $\mathbf{x}$ and (constant) physical variables $\mathbf{y}$.
-  • 7.
+  • 8.
     $\mathbf{x}$: A vector containing the available planning & control variables (i.e.
@@
 These variables are bounded such that $lb_{n}\leq x_{n}\leq ub_{n}$, where $lb_{n}$ is the lower bound, $ub_{n}$ is the upper bound.
-  • 8.
+  • 9.
     $\mathbf{y}$: A vector containing the physical variables (i.e. exogenous ’non-design’ variables) $y_{1},y_{2},...,y_{M}$.
-  • 9.
+  • 10.
     $g_{p}(O_{i}(\mathcal{N}(\mathbf{x},\mathbf{y}),F_{1,2,...,J}(\mathbf{x},\mathbf{y})),F_{1,2,...,J}(\mathbf{x},\mathbf{y}))$: Inequality constraint functions, which can be either objective function and/or performance function constraints.
-  • 10.
+  • 11.
     $h_{q}(O_{i}(\mathcal{N}(\mathbf{x},\mathbf{y}),F_{1,2,...,J}(\mathbf{x},\mathbf{y})),F_{1,2,...,J}(\mathbf{x},\mathbf{y}))$: Equality constraint functions, which can be either objective function and/or performance function constraints.
-  • 11.
+  • 12.
     $w^{\prime}_{k,i}$: Weights for each of the preference functions (sum of the number of preference functions equals $K*I$).
@@
 The preference aggregation is performed using the principles of the PFM theory (Barzilai, 2022).
-It is an integral part of the integrative preference maximisation method (IMAP), see Wolfert (2023).
+It is an integral part of the Integrative Preference Aggregation Method (IMAP), see Wolfert (2023).
+A best-fit solution is identified as a feasible plan that yields the highest aggregated preference.
+Note that individual preferences reflect the level of interest in specific objectives.
 To arrive at an aggregated preference for individual objectives, the IMAP method employs the mathematical operator, A, which determines the aggregated preference score using an algorithm called the A-fine Aggregator (see A).
@@
 The preference functions describe the relation of a stakeholders satisfaction towards a certain objective on a defined scale.
+In other words, preference — the only property of relevance in decision theory and equivalent to value — provides goal orientation to the objective function.
+Note: objective functions, in themselves, merely describe an objective and do not inherently imply a minimisation or maximisation direction — a distinction that sets IMAP apart from all classical multi-objective optimisation methods.
 The elicitation of the preference functions and associated individual weights is essential to reflect the relationship of the stakeholders interest towards a certain objective.
@@
 However, for a pure project control optimisation (no planning variables), the variables $\mathbf{x}$ are redefined as $\mathbf{x}=\mathbf{a}\mathbf{c_{i}}$, that is the product of the allocation of a control variables vector $\mathbf{a}$ (which takes binary value $a_{n}\in\{0,1\}$, where 0 reflects not allocated, or 1 reflects allocated) and the control impact vector $\mathbf{c_{i}}$ towards objective $i$, containing the impact of control variables $c_{1},c_{2},...,c_{N}$ directly related towards control variables $x_{1},x_{2},...,x_{N}$.
+Note that $x_{1..N}\in\{0,c_{1..N}\}$, which implies that the solution space is bounded.
 3 Odycon’s stochastic simulation & optimisation methodology
@@
 In case of a delay, the optimisation is initiated.
-(6) The optimisation is run by defining the objective functions, the preference functions and their related weights, in accordance with the prior established mathematical formulation for the optimisation (see section 2).
+(6) The optimisation is run by defining the objective functions, the preference functions (which defines the desired outcomes for the objective functions) and their related weights, in accordance with the prior established mathematical formulation for the optimisation (see section 2).
 Once defined, the optimisation is performed utilising the IMAP method.
@@
 The optimisation is executed within every iteration of the probabilistic MCS, aiming to select the most effective set of project planning variables and or control measures given the specific realisation of the network within that iteration.
+The
 The essence of the IMAP method lies in combining preferences aggregation (reflecting associative decision-making) with preference maximisation (reflecting human goal-oriented behaviour).
@@
 To read these examples here independently, the project descriptions are partly repeated.
-4.1 Strategic planning application
+4.1 Strategic Planning application : Agree First & Act Feasibly
 Technical context: Offshore floating wind (OFW) technologies are considered a solution for wind energy production in deep waters and regions with stronger wind velocities (Spring, 2020).
@@
 $\mathcal{N}(x_{1},x_{2},x_{3},\mathbf{y})$ (4)
-where the physical variables $\mathbf{y}$ contain the information on the respective vessel deck space.
+where the variables $\mathbf{y}$ contain the information on the respective vessel deck space or are related to the risks (and duration of the activity).
 To ensure that the project is executed by at least one vessel, the sum of all vessels on the project must be greater than one.
 The technical feasibility is guaranteed by a constraint ensuring the anchor’s resistance meets or exceeds the applied forces defined by further performance functions ($F_{4}$ ,…, $F_{7}$).
-For a thorough exposition of these formulations and their underlying principles, the reader is referred to Wolfert (2023) or Van Heukelum et al.
+For a complete exposition of these formulations and their underlying principles, the reader is referred to Van Heukelum et al.
 (2024).
@@
 The optimisation framework considers the following four objectives that form the link between the network performance function and the preference functions.
-For a thorough exposition of these formulations and their underlying principles, the reader is referred to Wolfert (2023) or Van Heukelum et al.
+For a complete exposition of these formulations and their underlying principles, the reader is referred to Wolfert (2023) or Van Heukelum et al.
 (2024).
@@
 Note, stakeholders can also reflect interest to all four objectives while the global weight distribution is constant.
-Stakeholder $k$ $w^{\prime}_{k,PD}$ $w^{\prime}_{k,S}$ $w^{\prime}_{k,C}$ $w^{\prime}_{k,PD}$ $w_{k}$
+Stakeholder $k$ $w^{\prime}_{k,PD}$ $w^{\prime}_{k,S}$ $w^{\prime}_{k,C}$ $w^{\prime}_{k,F}$ $w_{k}$
 Energy provider 0.30 0.20 0.00 0.00 0.50
@@
 Figure 3: Criticality index of number of small OCV’s (top left), large OCV’s (top right), barges (bottom left) and vessel combinations (bottom right) for both optimisation cases.
-Note, only combinations that stored during the simulation & optimization are considered in this figure.
-4.2 Dynamic control application
+Note, only combinations that stored during the simulation & optimisation are considered in this figure.
+4.2 Dynamic Control application: Adapt Flexibly
 Technical context: The Schiphol-Amsterdam-Almere (SAA) Project is the largest road construction project in the Netherlands in the period from 2012 to 2024.
@@
 First, the complex reality of project management is addressed by identification of the most effective and desired set of project planning & control variables given the uncertainties and stochastic factors while considering overall group satisfaction.
-The interplay with stakeholder-oriented concurrent objectives is addressed by the integration of the ’associative preference domain’, enabling an optimal solution that best fits the common purpose, using a-priori optimization.
+The interplay with stakeholder-oriented concurrent objectives is addressed by the integration of the ’associative preference domain’, enabling an optimal solution that best fits the common purpose, using a-priori optimisation.
 This approach supports satisfactory project planning and control, allowing stakeholders to meet their individual needs through collective performance.
@@
 Note, for preference aggregation as part of the optimisation, the following algorithm is used: https://github.com/Boskalis-python/a-fine-aggregator.
+Acknowledgements
+The authors would like to thank Dr. Ruud Binnekamp of Delft University of Technology for his valuable reflections and guidance in completing this paper.
 Appendix A Preference score aggregation

### v3

@@
 Abstract
-Strategic project planning and dynamic control are essential to ensure that complex projects are executed in a best-fit-for-common-purpose manner.
-When such strategies become inconceivable for humans, effective and efficient computer-aided decision support becomes crucial.
-Standard simulation-driven evaluation and a-posteriori decision-making are often used, rather than a combined simulation and optimisation approach that a-priori integrates the stakeholder interests and risk management.
-However, current planning and control methodologies often lack stochastic representation and/or associative multi-objective optimisation approaches within socio-technical reach.
-This paper introduces Odycon (Open design and dynamic control), a new purpose-driven project management by design methodology that resolves these shortcomings.
-A generic mathematical project planning and control statement is developed to integrate and represent both stakeholders preferences (human domain) and project system performances (technical domain) for effective project planning and dynamic control.
-Odycon integrates Monte-Carlo simulation (MCS) with the Integrative Maximisation of Aggregated Preferences (IMAP) optimisation method to develop a best-fit strategic plan and dynamic control measures.
-Its value is demonstrated through applications in offshore wind installation and a highway infrastructure projects, showcasing advances in associative design and decision-making, and aiming for a best-fit-for-common purpose synthesis across different complex project phases.
+Strategic project planning and dynamic control are essential to ensure that complex projects are both prepared and executed best-fit-for-common-purpose, guided by three interrelated strategies: (1) Agreeing First, (2) Acting Feasibly, and (3) Adapting Flexibly.
+When these strategies become too complex for humans to fully conceive and manage, effective computer-aided decision support becomes crucial.
+However, standard simulation-driven evaluation and a-posteriori decision-making are typically single-sided and technically focused focus, rather than applying a combined simulation-and-optimisation approach that a-priori integrates stakeholder interests and their mitigation behavior.
+Moreover, current planning and control methodologies often lack robust stochastic representations and associative multi-objective optimisation methods that capture the full socio-technical complexity while maximizing the potential within reach.
+This paper introduces Odycon (Open Design and Dynamic Control), a new purpose-driven project management methodology that provides an actionable solution to these challenges.
+It presents a generic mathematical framework for project planning and control that integrates stakeholder preferences (human domain) with system performances (physical domain), enabling more effective planning and dynamic control.
+Odycon integrates standard Monte Carlo simulation (MCS) with the novel Integrative Maximisation of Aggregated Preferences (IMAP) optimisation method to develop a best-fit strategic plan and the most effective mitigation control strategies.
+Its value is demonstrated through applications in offshore wind installation and highway infrastructure projects, showcasing advances in associative design and decision-making, and aiming for a best-fit-for-common-purpose synthesis across different complex project phases.
 keywords:
-Project management , Project planning & scheduling , Stochastic Project control , Risk management , Multi-objective optimisation , Preference & Performance-based optimization , Monte Carlo simulation , Preference function modelling
+Project management by Design , Human-centred Project planning & Dynamic Control , Risk management , Multi-objective optimisation , Preference-based optimization , Monte Carlo simulation , Preference function modelling
+journal: Journal of Construction Engineering and Management
 […]
@@
 And even when they start right, they can still lose direction.
-Advanced design and project management practices may still hit impasses because project managers often rely on retrospective analyses (such as those by Flyvbjerg (2024) or others like Yang et al.
-(2022)) and traditional decision-support methods, which offer little guidance for a constructive way forward.
-These challenges become particularly acute when decision-makers fail to recognize that their problems are part of a larger, interconnected whole — and that mutual interconnection and collaboration are key to unlocking complex projects that achieve a best-fit for a common purpose.
-This calls for moving beyond the traditional notion of fitness for purpose (total quality), which emphasizes meeting predefined requirements and delivering technically effective results.
-Instead, we must adopt a broader, value-driven perspective — one that focuses on what is socially desirable and physically feasible from a shared purpose.
+Advanced design and project management practices may still hit impasses because project managers often rely on retrospective analyses - such as those by Flyvbjerg (2024) or others like Yang et al.
+(2022) - and traditional decision-support methods, which offer little to no actionable guidance for a constructive, best-for-project way forward.
+These challenges become particularly acute when decision-makers fail to recognize that their problems are part of a larger, interconnected whole, and that mutual interconnection and collaboration are key to unlocking complex projects that achieve best-fit for a common purpose.
+This calls for moving beyond the traditional notion of fitness for purpose — often associated with total quality bound by scope, cost, and time (also known as project management triangle or triple constraint) and the delivery of technically effective results that meet predefined requirements.
+In today’s complex construction environments, this narrow, end-product-oriented view is no longer sufficient.
+Instead, we must adopt a broader, value-driven perspective that focuses on what is socially desirable and physically feasible guided by a common purpose from the very start and continuing throughout to the end.
 Best-fit for common purpose integrates stakeholder interests, dynamic context, and collective value creation within socio-physical reach.
@@
 However, as project size and uncertainty increase, initial estimates regarding scope, cost, and timelines are often exposed to inherent risks and uncertainties that cannot be entirely eliminated.
-Project success is therefore achieved through dynamic adaptation to emerging challenges—often not specified in a static contract—and through collaborative engagement among stakeholders to develop associative solutions.
+Project success is therefore achieved through dynamic adaptation to emerging challenges—often not specified in static contracts—and through collaborative engagement among stakeholders or decision-makers to co-develop associative solutions.
 Focusing solely on the final scope as defined in the contract, insisting on creating a fixed plan and adhering rigidly to it, is ultimately naïve and pseudo-realistic in the face of real-world project dynamics.
@@
 By integrating diverse stakeholder interests and accounting for project capabilities in the decision-making process, a more holistic outcome—referred to as the best-fit for common purpose—can be achieved, rather than settling for suboptimal compromises (Zhilyaev et al., 2022).
-This best-fit design emerges from the socio-technical interplay between stakeholder preferences (the human domain—what they want) and system performance capabilities (the technical domain—what it can deliver), as developed and demonstrated by Van Heukelum et al.
+This best-fit design emerges from the socio-technical interplay between stakeholder preferences (the human domain—what they want) and system performance capabilities (the physical domain—what it can deliver), as developed and demonstrated by Van Heukelum et al.
 (2024).
@@
 In this paper, all the aforementioned considerations are taken into account in the development of a new open-design-driven dynamic project planning and control methodology, named Odycon (acronym: Open Design and Dynamic Control).
-Odycon builds on the core principles and optimization method of the Open Design Systems (Odesys) methodology (Wolfert, 2023).
+Odycon builds on the core principles and design optimization method of the Open Design Systems (Odesys) methodology (Wolfert, 2023).
 The Odesys integrative maximisation of aggregated preferences (IMAP) optimization method has proven effective in achieving best-fit-for-common-purpose design solutions, where all stakeholder interests are accounted for and translated into a common preference domain to identify an overall group optimum.
-Notably, Odesys/IMAP is a novel preference-based optimization methodology where, instead of directly minimizing or maximizing individual objectives, the rigorous mathematical optimization focuses solely on maximizing the aggregated preference of stakeholder objectives within physical reach.
+Notably, Odesys/IMAP is a novel preference-based optimization methodology where, instead of directly minimizing or maximizing individual objectives, the rigorous mathematical optimization focuses solely on maximizing the aggregated preference of stakeholder objectives within physical reach while removing fundamental modelling errors of standard multi-objective optimization methods (Van Heukelum et al., 2024).
 So far, the Odesys/IMAP method enables iterative group design and decision-making in a deterministic manner.
@@
 (2022, 2021), will be enhanced from a single-cost objective orientation to optimally facilitate stochastic, multi-objective project control.
-In doing so, Odycon constitutes a truly novel project management-by-design methodology.
-To summarize, the following Odycon advances are presented in this paper: 1.
+In doing so, Odycon constitutes a truly novel human-centred project management-by-design methodology.
+To summarize, the following Odycon advances are presented in this paper:
+  1. 1.
     A general framework for integrated strategic planning and dynamic control, combining probabilistic project schedule simulations with preference-based multi-objective design optimisation.
 This enables dynamic adaptation to project-specific conditions and stakeholder objectives, thereby enhancing associative decision-making in complex construction environments.
-2.
+  2. 2.
     A direct linkage between project feasibility, project management objectives, and stakeholder decisions, resulting in a unified integration of network performance, stakeholder values, and associative planning and control processes.
-3.
+  3. 3.
+    A human-centred, actionable project management decision-support model that overcomes fundamental modelling errors in classical multi-objective optimisation and Monte Carlo planning by fully integrating human preferences with physical performances, eliminating invalid aggregation and Pareto-front assumptions, and accounting for human-oriented mitigation behavior (Kammouh et al., 2021; Van Heukelum et al., 2024).
+  4. 4.
     A pure preference-based aggregation mechanism for multi-objective decision-making in project management.
-This translates diverse objectives and constraints into a shared preference domain, enabling an optimised group synthesis without relying on monetisation.
+This translates diverse objectives and constraints into a shared preference domain, enabling an optimised group synthesis without relying on monetisation (Wolfert, 2023).
 Rather than selecting the lowest cost solution, the approach supports identifying the solution with the highest combined value for both project outcomes and people—that is, a best-fit for common purpose plan and set of control measures.
-4.
+  5. 5.
     A methodological advancement from PII to PIII in project management practice.
@@
 Next, the underlying framework and core concepts of the stochastic simulation and optimisation methodology are explained, including a detailed description of the method’s design, within which the IMAP optimisation is embedded.
-The methodology is first demonstrated using a strategic planning case (with no control variables) involving an offshore wind installation project.
-In this scenario, a service provider and a marine contractor act as the primary decision-makers and are able to agree on a shared scope, resulting in an actionable and feasible optimal plan.
-Following this, a dynamic control case (with no planning variables) is presented, applied to a highway infrastructure project.
-This demonstrator features associative mitigation control by both a contractor and a highway agency, illustrating how to arrive at a best set of control measures to enable flexible adaptation across two distinct project management phases.
+The methodology is first demonstrated using a strategic planning case - Agreeing First & Acting Feasibly (with no control variables) - involving an offshore wind installation project, where a service provider and a marine contractor are the key decision-makers.
+This is followed by a dynamic control case - Adapting Flexibly (with no planning variables) - applied to a highway infrastructure project.
+This demonstrator features associative mitigation control by both a contractor and a highway agency, illustrating how to identify a best set of control measures to enable flexible adaptation across two distinct project management phases.
 The paper concludes with recommendations for further development and reflects on the added value of the newly developed stochastic simulation and optimisation methodology.
@@
   • 1.
-    $U$: Utility function that needs to be maximized for a best-fit configuration, is addressed using the Genetic Algorithm software tool, preferendus (Van Heukelum et al., 2024).
+    $U$: Utility function that needs to be maximized for a best-fit configuration, is addressed using a Genetic Algorithm software tool, preferendus (Van Heukelum et al., 2024).
   • 2.
-    $A$: Aggregated preference score determined as part of the IMAP optimisation method (see A).
+    AA: Aggregated preference score determined as part of the IMAP optimisation method (see A).
   • 3.
@@
   • 6.
-    $\mathcal{N}(\mathbf{x},\mathbf{y})$: Performance function that describes the project network, depending on planning and/or control variables $\mathbf{x}$ and physical variables $\mathbf{y}$.
-Here, $\mathcal{N}$ is a parametric graph representation of a project planning network, composed of nodes (activities) $V$ and edges (logical links) $E(\mathbf{x},\mathbf{y})$.
-Each node $v\in V$ has attributes such as duration $D_{v}(\mathbf{x},\mathbf{y})$, risk $R_{v}(\mathbf{x},\mathbf{y})$, and cost $C_{v}(\mathbf{x},\mathbf{y})$, while each edge $e\in E(\mathbf{x},\mathbf{y})$ has attributes such as weight $W_{e}(\mathbf{x},\mathbf{y})$ and capacity $C_{e}(\mathbf{x},\mathbf{y})$.
+    $F_{1,2,...,J}(\mathbf{x},\mathbf{y})$: Performance functions that describe the object behaviour, depending on the planning and/or control variables $\mathbf{x}$ and (constant) physical variables $\mathbf{y}$.
   • 7.
-    $F_{1,2,...,J}(\mathbf{x},\mathbf{y})$: Performance functions that describe the object behaviour, depending on the planning and/or control variables $\mathbf{x}$ and (constant) physical variables $\mathbf{y}$.
-  • 8.
     $\mathbf{x}$: A vector containing the available planning & control variables (i.e.
@@
 These variables are bounded such that $lb_{n}\leq x_{n}\leq ub_{n}$, where $lb_{n}$ is the lower bound, $ub_{n}$ is the upper bound.
+  • 8.
+    $\mathbf{y}$: A vector containing the physical variables (i.e. exogenous ’non-design’ variables) $y_{1},y_{2},...,y_{M}$.
   • 9.
-    $\mathbf{y}$: A vector containing the physical variables (i.e. exogenous ’non-design’ variables) $y_{1},y_{2},...,y_{M}$.
+    $g_{p}(O_{i}(\mathcal{N}(\mathbf{x},\mathbf{y}),F_{1,2,...,J}(\mathbf{x},\mathbf{y})),F_{1,2,...,J}(\mathbf{x},\mathbf{y}))$: Inequality constraint functions, which can be either objective function and/or performance function constraints.
   • 10.
-    $g_{p}(O_{i}(\mathcal{N}(\mathbf{x},\mathbf{y}),F_{1,2,...,J}(\mathbf{x},\mathbf{y})),F_{1,2,...,J}(\mathbf{x},\mathbf{y}))$: Inequality constraint functions, which can be either objective function and/or performance function constraints.
+    $h_{q}(O_{i}(\mathcal{N}(\mathbf{x},\mathbf{y}),F_{1,2,...,J}(\mathbf{x},\mathbf{y})),F_{1,2,...,J}(\mathbf{x},\mathbf{y}))$: Equality constraint functions, which can be either objective function and/or performance function constraints.
   • 11.
-    $h_{q}(O_{i}(\mathcal{N}(\mathbf{x},\mathbf{y}),F_{1,2,...,J}(\mathbf{x},\mathbf{y})),F_{1,2,...,J}(\mathbf{x},\mathbf{y}))$: Equality constraint functions, which can be either objective function and/or performance function constraints.
-  • 12.
-    $w^{\prime}_{k,i}$: Weights for each of the preference functions (sum of the number of preference functions equals $K*I$).
+    $w^{\prime}_{k,i}$: Weights for each of the preference functions (sum of the number of preference functions equals KIK*I).
 The global weights for the relative importance of stakeholders is defined as $w_{k}$.
@@
 Note that individual preferences reflect the level of interest in specific objectives.
-To arrive at an aggregated preference for individual objectives, the IMAP method employs the mathematical operator, A, which determines the aggregated preference score using an algorithm called the A-fine Aggregator (see A).
+To arrive at an aggregated preference for individual objectives, the IMAP method employs the mathematical operator A, which computes the aggregated preference score using an algorithm called the A-fine Aggregator (see A).
 Remark 2: Preference functions
@@
 Variables Description Boundaries Anchor deck space Day rate $R_{i}$ utilisation prob. $p_{i}$ CO2 emissions $E_{i}$
-$F1=x1$ small OCV $0\leq x1\leq 3$ 8 $47k\text{\texteuro}$ 0.7 30 $30\text{\ }t/day$
-$F2=x2$ large OCV $0\leq x1\leq 2$ 12 $55k\text{\texteuro}$ 0.8 40 $40\text{\ }t/day$
-$F3=x3$ barge $0\leq x1\leq 2$ 16 $35k\text{\texteuro}$ 0.5 35 $35\text{\ }t/day$
+$F1=x1$ small OCV $0\leq x1\leq 3$ 8 $47k\text{\texteuro}$ 0.7 $30\text{\ }t/day$
+$F2=x2$ large OCV $0\leq x1\leq 2$ 12 $55k\text{\texteuro}$ 0.8 $40\text{\ }t/day$
+$F3=x3$ barge $0\leq x1\leq 2$ 16 $35k\text{\texteuro}$ 0.5 $35\text{\ }t/day$
 Project activities
@@
 Table 3: Strategic planning application: Risk events.
-Risk Event Risk duration [$days$] Affected Activity Probability $p_{r}$
+Risk Event Risk duration [daysdays] Affected Activity Probability $p_{r}$
 $a_{i}$ $m_{i}$ $b_{i}$
@@
 The following objective functions are rewritten based upon to suit the needs of the Odycon framework.
-  1. 1.
-    Project duration: the project duration $\Delta$ extracted from the network performance function expressed as:
+Objective 1: Project duration
+The project duration $\Delta$ extracted from the network performance function expressed as:
 $O_{1}=O_{PD}=\Delta(\mathcal{N}(x_{1},x_{2},x_{3},\mathbf{y}))$ (5)
-  2. 2.
-    Installation cost: the project cost depends on the day rates of the vessels $R_{i}$ multiplied with their active duration on the project $t_{i}$ given the network performance function, and the cost of the installed anchors, expressed as:
+Objective 2: Installation cost
+The project cost depends on the day rates of the vessels $R_{i}$ multiplied with their active duration on the project $t_{i}$ given the network performance function, and the cost of the installed anchors, expressed as:
 $O_{2}=O_{C}=(815M_{a}+40,000)n_{a}+\sum_{i=1}^{3}x_{i}R_{i}t_{i}(\mathcal{N}(x_{1},x_{2},x_{3},\mathbf{y}))$ (6)
-  3. 3.
-    Fleet utilisation: utilisation is represented by the probability of a vessel being better utilised in another project and is defined as:
+Objective 3: Fleet utilisation
+Fleet utilisation is represented by the probability of a vessel being better utilised in another project and is defined as:
 $O_{3}=O_{F}=\prod\limits_{i=1}^{3}p_{i}^{x_{i}}$ (7)
-  4. 4.
-    Sustainability (CO2 emissions): emissions are defined as the daily vessel emissions $E_{i}$ multiplied with their active duration on the project $t_{i}$ given the network performance function:
+Objective 4: Sustainability (CO2 emissions)
+Emissions are defined as the daily vessel emissions $E_{i}$ multiplied with their active duration on the project $t_{i}$ given the network performance function:
 $O_{4}=O_{S}=\sum_{i=1}^{3}x_{i}E_{i}t_{i}(\mathcal{N}(x_{1},x_{2},x_{3},\mathbf{y}))$ (8)
@@
 The numerical results can be found in Table 5.
-Table 5: Evaluation of cases with 2000 MCS iterations ($O_{PD}$ in days, $O_{C}$ in $10^{7}$, $O_{F}$ in number of vessels [-], $O_{S}$ in tonnes).
+Table 5: Evaluation of cases with 2000 MCS iterations ($O_{PD}$ in days, $O_{C}$ in $10^{7}$ €, $O_{F}$ in number of vessels [-], $O_{S}$ in tonnes).
 Where the MOO case considers $w^{\prime}_{1,PD}=0.30,w^{\prime}_{1,S}=0.20,w^{\prime}_{2,C}=0.35,w^{\prime}_{2,F}=0.15$.
@@
 The optimisation framework considers the following four objectives that form the link between the network performance function and the preference functions.
-Objective project duration
+Objective 1: project duration
 The project duration $\Delta$ is extracted from the network performance function and can be expressed as follows:
@@
 where $O_{PD}$ is expressed in days.
-Objective project control cost
+Objective 2: project control cost
 The project control cost is defined as the sum of the cost impact of active control measures.
@@
 Note, the project duration after implementation of control measures $\Delta$ is extracted from the network performance function $\mathcal{N}(a_{n}*c_{2,n},\mathbf{y})$, see Equation 9.
-Objective traffic nuisance
+Objective 3: traffic nuisance
 During construction, the experience of a road user (considered as traffic nuisance) is heavily impacted.
@@
 However, this adaptation cannot be visualised.
-Table 7: Evaluation of cases with 2000 MCS iterations ($O_{PD}$ in days, $O_{C}$ in $10^{6}$, $O_{N}$ in nuisance level [0,10]).
+Table 7: Evaluation of cases with 2000 MCS iterations ($O_{PD}$ in days, $O_{C}$ in $10^{6}$ €, $O_{N}$ in nuisance level [0,10]).
 Where the MOO case considers $w^{\prime}_{1,PD}=0.70,w^{\prime}_{2,C}=0.15,w^{\prime}_{1,N}=0.15$.
@@
 This is expressed mathematically as follows:
-Minimise $\textit{Minimise}\text{\ }S_{i}=\sum_{j=1}^{J}w_{j}*(z_{i,j}-P_{i}^{*})^{2}$ (16)
+$\textit{Minimise}\text{\ }S_{i}=\sum_{j=1}^{J}w_{j}*(z_{i,j}-P_{i}^{*})^{2}$ (16)
 Note that since the search is for a single representative aggregated preference score, the model function $f(x_{i,j},\beta_{i})$ from the classical weighted least square method is replaced by $P_{i}^{*}$.